### v1

# Multimodal Banking Dataset: Understanding Client Needs through Event Sequences

Abstract
Financial organizations collect a huge amount of data about clients that typically has a temporal (sequential) structure and is collected from various sources (modalities). Due to privacy issues, there are no large-scale open-source multimodal datasets of event sequences, which significantly limits the research in this area. In this paper, we present the industrial-scale publicly available multimodal banking dataset, MBD, that contains more than 1.5M corporate clients with several modalities: 950M bank transactions, 1B geo position events, 5M embeddings of dialogues with technical support and monthly aggregated purchases of four bank’s products. All entries are properly anonymized from real proprietary bank data. Using this dataset, we introduce a novel benchmark with two business tasks: campaigning (purchase prediction in the next month) and matching of clients. We provide numerical results that demonstrate the superiority of our multi-modal baselines over single-modal techniques for each task. As a result, the proposed dataset can open new perspectives and facilitate the future development of practically important large-scale multimodal algorithms for event sequences.
HuggingFace Link: https://huggingface.co/datasets/ai-lab/MBD
Github Link: https://github.com/Dzhambo/MBD
1 Introduction
The ever-growing digital footprint of consumers has led to an explosion of data collected by businesses across all industries. Financial institutions, in particular, gather a vast amount of information about their clients over time, encompassing various aspects of their financial activities. This data, spanning extended periods, is typically annotated with temporal information, forming what is known as event sequences [8, 14, 23, 37]. Description of an event usually contains several heterogeneous fields, numerical and categorical, and processing of event sequence is a separate problem in machine learning different from techniques applied to time series or texts.
An important property of event sequences is that these data are often gathered from multiple sources or channels, rendering multimodal. The comprehensive analysis of such data in a joint way is critical for informed decision-making processes within enterprises. The scarcity of publicly available datasets in this domain hampers research and development efforts. Indeed, the significant progress in multimodal data for natural language processing, computer vision, audio, etc., is due to the appearance of large-scale datasets, such as LAION, DataComp, etc. [16]. Unfortunately, though several datasets of event sequences are used in research, e.g., credit card transactions [30] or MIMIC [21], they are either small or contain only one modality. Thus, a gap exists in methodologies designed to tackle the complexity of multimodal event sequence data.
To bridge this gap, this paper introduces the Multimodal Banking Dataset (MBD), an unprecedented open-source resource encompassing extensive multichannel event sequence data of banking corporate clients. This dataset is the largest of its kind, featuring detailed records of approximately 1.5 million clients across four distinct modalities: money transfers (about 950 million events), geo position data (around 1 billion events), technical support dialog embeddings (approximately 5 million entries), and monthly aggregated bank product purchases categorized into four types. Each modality encompasses roughly one or two year of historical, time-annotated data, making it a rich resource for analyzing the dynamics of client behavior over time.
The MBD dataset enables the exploration of several critical business problems in a multimodal context. Specifically, it facilitates research in such tasks as purchase prediction for the following month (campaigning), and matching of different modalities for the same clients. These problems can all be addressed by leveraging the temporal and spatial attributes inherent in the dataset. Moreover, we provide benchmarks for these tasks using the MBD. Importantly, all the data are anonymized to protect client privacy, and our research confirms that the anonymization process preserves the consistency of model performance between original and anonymized data. This paper underscores the potential of MBD as a foundational tool for advancing research in multimodal sequential data analysis, fostering innovation, and enhancing decision-making processes in the financial sector.
2 Dataset
Modern innovative banking institutions actively develop AI technologies for customizing their human-oriented technologies and making everyday decisions. A superior level of technologies will lead to new cases of customers’ experience, which should form a competitive advantage of services provided, including the speed, accuracy, and price of customer services, including personal credit conditions, individual finance strategy, etc.
One of the main benefits of using AI is the ability to analyze large amounts of customer data. This helps banks better understand their customers’ needs and offer them the most suitable products and services. Additionally, AI can protect customers from fraud and prevent financial losses. More accurate forecasting of financial risks associated with lending or investing allows us to provide more favorable conditions to more reliable clients.
This strategy is more effective if there is a lot of data, so banks strive to accumulate as much information as possible.
Being a team of a bank that stores petabytes of data about bank processes and clients, we understand the urgent needs of fintech data scientists in large sets of publicly available temporal data from various sources (bank transactions, client locations, purchased products, etc.) to drive innovation and scientific discovery. Unfortunately, the number of appropriate datasets is limited because providing such data comes with certain risks. Typically, banks are wary of sharing their data due to potential leaks of confidential information or violations of data protection regulations. In addition, the data is considered to have commercial value, and companies do not want to disclose it. To mitigate these risks, it is required to take all necessary steps to ensure data security and remove any identifying information. This allows information to be shared without violating client confidentiality, but this procedure requires significant effort from engineers, managers, and lawyers. Though there exist a small number of properly anonymized banking datasets, such as credit card transactions [30] or AlphaBattle [2], to the best of our knowledge, there are no publicly available large multimodal temporal datasets for banks.
Thus, in this paper, we introduced the first large-scale multimodal banking dataset to support future research on multimodal techniques for event sequences. In particular, we selected several practically important tasks, such as campaigning, i.e., prediction if a client would purchase some of four rather popular products in the next month. Each client is described by sequences of typical bank data: transactions, geo positions where the customer used the bank application, and dialogues with technical support. These data sources ideally highlight the main difficulties in developing multimodal models, namely, asynchronous events in different modalities, various intensities of events, rare/irregular events, and even the absence of some modalities for many clients. Based on our dataset, the researchers will be able to fully take into account cross-modal connections of sequences from multiple sources at the level of individual events.
The dataset was collected using the following procedure. At first, we selected a complete sample of clients for two year (2021, 2022) to be able to cover all seasons. Among all customers who had the opportunity to purchase at least one of four products during 2022, we randomly selected 2,186,230 clients, among which 1M customers are labeled by monthly aggregated purchases of each of four products in each month. For these clients, we collected 947,899,612 bank transactions, 1,117,213,760 geo position events, and 5,080,781 dialogues with technical support. Our data raise typical practical challenges for training multimodal models. For example, many clients do not have all three modalities simultaneously because they can never make a transaction, call tech support, or leave their geo trace while running the bank application. Next, all the data were properly anonymized to guarantee the confidentiality and privacy of customer information. As a result, it becomes impossible to recover back to real clients from our anonymous data. We will show in the experimental study that such anonymization still let us extract valuable information about clients.
To demonstrate the complete transformation of data from various sources, Fig. 1 briefly shows example data for each modality, and the temporal structure of the data for a campaigning model is presented. Let us introduce further details for each modality in the following subsections.
2.1 Modalities
1. Bank transactional data are financial operations (events) carried out between different clients. The sequence of transactions can uniquely characterize the client [8], so this modality plays one of the most important roles in financial planning and recommendations. Thus, the main component of our MBD dataset is each client’s transactional history, represented by an event with a timestamp and various attributes of the anonymized counterparty. Clients have 638 transactions on average. Fig. 2 in Appendix A contains the detailed distribution of clients for different lengths of the transaction history. The list of attributes is shown in Table 4.
2. Dialogues. The dialogue data are transcriptions of customer calls to technical support and negotiations between the client and its manager. It is an extremely important source of information about client needs and problems [11]. The audio utterance is fed into a commercial Speech-to-Text algorithm. Personal information, e.g., the client’s name, is detected in the text and masked. To further anonymize the dialogue, we feed its text into a pre-trained NLP model111 https://huggingface.co/ai-forever/ruBert-base and save the resulting embeddings of size 768 in dialogue modality. An example of dialogue data is presented in Fig. 3 in Appendix A. Only 46% of customers contact support and have records of conversations, 98% of them have no more than 10 dialogues per year. Fig. 4 contains the detailed distribution of clients for various numbers of dialogues.
3. Geostream data contains a sequence of geo-coordinates of a client. To anonymize this modality, the coordinates are encoded using geohashes222https://pypi.org/project/pygeohash/, a geocoding system that converts a geographic location into a short string. Each unique geohash corresponds to a region on the Earth’s surface. It is possible to adjust the accuracy and size by removing characters from the end of the code. In our dataset, the coordinates were encoded with a precision of 4, 5, and 6 characters, representing cells of different sizes on the map. In our dataset, there are 43,999 distinct values of geohash_4, 347,698 numbers of geohash_5, and 2,264,404 most precise locations (geohash_6).
An example of geo data is presented in Fig. 5 in Appendix A. The clients have an average of 910 geo-points in the time period considered. Detailed information about the distribution of their length over the clients is available in Fig. 6.
4. Products purchases. Our dataset is a valuable resource for analyzing the bank customers’ needs. In particular, it contains information about purchases of four different banking products for each month of 2022. To predict a purchase in a certain month, it is necessary to take events (transactions, geo, dialogues) strictly before the beginning of this month. Therefore, the date range for the purchases dataset is shifted by 1 month, i.e., information is available from February 1, 2022, to January 31, 2023. The campaigning task is a multi-label classification problem, i.e., we store a binary label for each product that indicates whether it was purchased by a customer in a certain month. Example of data about purchases is shown in Fig. 8 (Appendix A). The peculiarity of this dataset is its imbalance, which is specific to this type of business task: 81% of clients have no purchases, 15% have one, and the remaining 4% have two or more purchases. A historical overview over 12 months allows us to model the customer behavior dynamic and predict the date of purchase more accurately.
2.2 Data anonymization
Our dataset contains no personal or confidential information whatsoever. Nevertheless, the event sequences are detailed enough that it could be possible to compare individuals from the publicly accessible portion of the dataset with the original proprietary data. To mitigate this risk, noise has been introduced to the data, ensuring that such comparisons and identification are impossible. The noise patterns were selected by our bank’s internal security department. These patterns are applied locally, preserving the overall structure of the data. The specific parameters of the noise are not disclosed to prevent potential attacks on the dataset.
All ID fields were hashed with a random salt. All categorical field values was mapped to enumerated indexes. Random noise was added to numerical fields. The noise was added to date fields with preserving hour of original date, which may be the cause the shuffle of local sequence. The dialogue embedding space was divided into regions, which were then shuffled.
3 Benchmark
In this paper, we formulate two practically important tasks using our new dataset: campaigning and client matching. This section contains details about baseline methods and testing protocols.
3.1 Methods
To establish performance baselines, we implement several widely-adopted architectures. Our approach prioritizes unsupervised and semi-supervised methods [10], enabling the training of a general-purpose encoder on unlabeled sequential data. Additionally, we incorporate supervised methods that allow for the immediate training of the encoder in a fully supervised manner.
The following techniques are implemented in our benchmark to extract features for Transactions and Geostream modalities:
1.
Aggregation Baseline that contains hand-crafted aggregation statistics [8]: 3,000 and 4 features in total for sequences of transactions and geodata, respectively.
2.
CoLES (Contrastive Learning for Event Sequences), a self-supervised contrastive model [8] specially developed to obtain representations of such event sequences as bank transactions. The sequence encoder is a GRU (Gated Recurrent Unit) with a hidden size 256.
3.
Two Tabular Transformers from IBM [30]. The first model, TabBERT, adapts BERT to event sequences such as bank transactions. The second model, TabGPT, was initially proposed to generate synthetic tabular sequences. Both models extract 256-dimensional embeddings of an input event sequence. After, we pool output embeddings of the client in result embedding of size 1024, calculating min, max, mean, and std.
To obtain representation of a sequence of dialogues, we implemented several straightforward techniques to aggregate the sequence of embeddings of each dialog of a client: 1) mean pooling of all embeddings; and 2) use only the most recent embedding for the date of interest.
To demonstrate the potential of multimodal processing of event sequences, we implemented traditional fusion techniques, namely, 1) blending, i.e., a weighted sum of scores (estimates of class posterior probabilities) from single-modal classifier, and 2) late fusion, i.e., learning a classifier for concatenated embeddings from all modalities [20]
For supervised methods, we use the following architectures: for unimodal data, we utilize an GRU encoder, while for multimodal data, we employ different GRU encoders for each modality and embeddings from all encoders are concatenated. The resulting neural network is trained using binary cross-entropy (BCE) loss in multilabel setting. For the GRU encoders we used hidden size 32.
3.2 Tasks
3.2.1 Campaigning task
In this task, it is required to predict the customer’s propensity to purchase four different products in the next month (Fig. 1) given sequences of transactions, geo locations, and dialogues from the beginning of this month. Solutions to this problem are used to plan marketing campaigns and prepare sales communications through various communication channels with the client.
We implement an out-of-fold validation protocol to conduct our experiments. The entire client dataset is divided into five folds: four folds are used for training, while the remaining fold is designated for testing. Both the training and testing sets are made publicly available alongside our dataset to enable future researchers to compare performance metrics.
The baseline methods outlined earlier are applied as follows. First, we train our models using the training set. Considering the temporal structure of our target, we compute the embedding of each client’s history up to one month, focusing on the presence of the target product (Fig. 1). We then evaluate the model using the multi-label classification metric ROC AUC across the 12 months of 2022 and for four binary product labels.
3.2.2 Multimodal Matching
Multimodal matching [41] involves aligning and comparing different modalities to identify meaningful relationships or connections between them. Frequently, data from multiple sources for the same client are matched using predefined rules or heuristics, which may not always yield optimal results. To enhance the accuracy of this process, specialized identification algorithms are required to more precisely compare different modalities.
For the matching task, we employed a framework analogous to CLIP [31]. We utilized GRU encoders to embed pairs of samples from two input modalities, labeling them as either positive matches (i.e., data from the same client) or negative matches (i.e., data from different clients). The model was trained using the InfoNCE loss function [13], which maximizes similarity for positive pairs while minimizing it for negative pairs. To assess the model’s performance, we used Recall@1, Recall@50 and Recall@100 metrics.
4 Experiments
Our models, experiments, and training procedures were implemented in Python, leveraging PyTorch and PyTorch Lightning for deep learning tasks, and PySpark for distributed data processing. We trained the neural networks using NVIDIA V100 GPU, while the boosting models were trained on computational clusters equipped with 600 cores. The reported experiments, including extensive hyperparameters optimization, required approximately 500 hours of computation time.
4.1 Model and training hyperparameters
We employed the unsupervised baseline methods (CoLES, TabGPT, TabBERT, Aggregation) with default hyperparameters from the pytorch-lifestream framework. The PyTorch implementation of the Adam optimizer was utilized, with an initial learning rate of 0.001, coupled with the StepLR scheduler. Models were saved based on the lowest validation loss or the highest validation unsupervised metrics [33], evaluated after each training epoch, with training of 15 epochs.
We employed 24-dimensional embeddings for categorical features and clipped the number of categories for features with many unique values. We applied either an identical mapping or a logarithmic transformation for numerical features. We used the gradient boosting algorithm available in PySpark ML for downstream tasks.
4.2 Campaigning task
In this Subsection, we provide experimental results for the campaigning task in unimodal and multimodal baselines.
One of the main objectives of our paper is to provide a real data benchmark to facilitate the development of multimodal algorithms. To achieve this, we must demonstrate that an algorithm outperforming another on our public benchmark will similarly outperform it on real data.
Table 1 shows that the transaction modality plays the most crucial role in achieving accurate classification. In contrast, dialogues and geostream, when used in a unimodal setting, performed only slightly better than a random estimator. However, as shown in Table 2, the predictive performance improved significantly in the multimodal setting through the integration of additional modalities. The overall trend indicates a consistent improvement in validation metrics as more modalities are incorporated. Specifically, the multimodal late fusion approach enhanced predictive accuracy by 1–2% when adding a single data source to the transaction stream, and by up to 3% when utilizing all available data sources.
Models trained on public datasets generally exhibit higher performance scores, which is an anticipated outcome of the data anonymization process. Nevertheless, the relative ranking of different methods remains consistent across both MBD and private datasets. Fig. 7 in Appendix B demonstrates a strong correlation in performance metrics between the public and private datasets, with a Kendall-tau correlation coefficient of 0.94.
Comprehensive results for each modality and all possible fusion combinations of multiple modalities are provided in Appendix B. We denote our modalities: geostream, transactions, and dialogues are represented as Geo, Trx, and Dialogs, respectively. We specify the method with the modality to indicate that a particular method was applied. For example, if embeddings from the CoLES model are used on transactions, we denote this as TrxCoLES.
| Modalities | Methods | MBD | Private dataset |
| Dialog | Mean | 0.595 0.002 | 0.604 0.001 |
| Last | 0.586 0.001 | 0.590 0.001 | |
| Supervised | 0.540 0.012 | 0.538 0.010 | |
| Geo | Aggregation | 0.555 0.001 | 0.554 0.001 |
| TabGPT | 0.621 0.003 | 0.622 0.001 | |
| TabBERT | 0.603 0.002 | 0.596 0.002 | |
| CoLES | 0.598 0.004 | 0.601 0.004 | |
| Supervised | 0.566 0.002 | 0.563 0.002 | |
| Trx | Aggregation | 0.783 0.002 | 0.791 0.017 |
| TabGPT | 0.802 0.002 | 0.796 0.000 | |
| TabBERT | 0.762 0.004 | 0.754 0.011 | |
| CoLES | 0.773 0.002 | 0.772 0.002 | |
| Supervised | 0.819 0.002 | 0.813 0.002 |
| Methods | Modalities | MBD | Private dataset |
| CoLES | Trx | 0.773 0.002 | 0.772 0.002 |
| Trx + Geo | 0.775 0.002 | 0.772 0.002 | |
| Trx + Dialog | 0.781 0.001 | 0.776 0.002 | |
| Trx + Dialog + Geo | 0.783 0.001 | 0.777 0.002 | |
| TabGPT | Trx | 0.802 0.001 | 0.796 0.000 |
| Trx + Geo | 0.800 0.001 | 0.796 0.001 | |
| Trx + Dialog | 0.810 0.002 | 0.805 0.001 | |
| Trx + Dialog + Geo | 0.808 0.001 | 0.802 0.001 | |
| Supervised | Trx | 0.819 0.002 | 0.813 0.002 |
| Trx + Geo | 0.819 0.001 | 0.815 0.001 | |
| Trx + Dialog | 0.821 0.0006 | 0.817 0.0006 | |
| Trx + Dialog + Geo | 0.824 0.001 | 0.821 0.001 | |
| TabBERT | Trx | 0.762 0.004 | 0.754 0.011 |
| Trx + Geo | 0.764 0.004 | 0.756 0.011 | |
| Trx + Dialog | 0.773 0.003 | 0.765 0.009 | |
| Trx + Dialog + Geo | 0.775 0.003 | 0.766 0.009 |
4.3 Multimodal matching task
In this subsection, we present the results of our proposed multimodal matching benchmark, summarized in Table 3, which includes both transactions and dialogues. Detailed results for other modalities can be found in Appendix B, in Table 9. In this evaluation, we report recall@1, recall@50, and recall@100 for each modality, measured in both directions. For example, in the case of the transaction and geostream pair, we compute both Trx2Geo and Geo2Trx, allowing for an evaluation of alignment from both perspectives.
Our analysis reveals considerable variation in performance across different modality pairs. Specifically, dialogue data consistently exhibits weaker matching performance compared to other modalities, such as transactions and geostream, which demonstrate significantly stronger alignment. This disparity suggests potential limitations within the dialogue modality, indicating that it may offer less complementary or aligned information. Alternatively, the unique structure of dialogue data may necessitate more sophisticated or specialized techniques for effective integration with other modalities.
These findings provide important insights for the advancement of multimodal matching methods, underscoring the need for further refinement.
| Recall@1 | Recall@50 | Recall@100 | |
| Trx2Geo | 0.006 0.0003 | 0.1967 0.002 | 0.303 0.004 |
| Geo2Trx | 0.004 0.0003 | 0.1624 0.002 | 0.262 0.004 |
5 Related Work
5.1 Financial data
The multitude of services and processes in banks generates a variety of data that can be considered as modalities. Early works such as [29, 26] use feature processing techniques that remove multimodal complexity and present data in tabular form. The Amex dataset[3] improves information content and complexity. Here, a wide range of different financial aggregates are presented as a sequence of historical slices. The development of deep learning methods has led to the ability to work with complexly structured data, such as sequences of events. Quite a few datasets [1, 5, 2], mostly monomodal, presented mainly at ML competitions. To work with such data, both supervised[7, 9, 36] and unsupervised methods[30, 8, 32] are used. Multimodal financial sequential data proposed in DataFusion2022 [4]. There are two sequential modalities, transaction and web clickstream, and two downstream tasks: matching and education level prediction. However, this is an extremely small dataset of 17K clients, and no accurate baseline model is available.
5.2 Other event sequence domains
Temporal point process[28, 40] model streams of discrete events in continuous time by constructing a neurally multivariate point process. The authors use a large collection of datasets from different types of modalities: Media (Retweets[39], MemeTrack[24], Amazon[6], IPTV[25]), Medical (MIMIC-II[22]), Social(Stack Overflow[24], Linkedin[38]) and Financial (Transaction[15]) data. All datasets are independent, and each one is single-modal. EventStreamGPT [27] uses multimodal medical record datasets, MIMIC-IV[21]. The authors propose a GPT-like approach for continuous-time event sequences. The structure of this dataset is close to the MDB. But it must be emphasized that financial data, unlike medical data, contains longer chains of events, more regular patterns, and individual transactions are less informative. These features may influence the quality of the methods used.
5.3 Geostream and dialogues
Geodata is used for various tasks. One of the uses of geo is a visualization of analytics on a map [17]. However, geo is used here not as a separate modality but as additional tags to the mainstream of tweets. Mobile marketers[12] use geo-targeting for pricing and send personalized recommendations. In [35] geo hashes are used for user mobility detection and prediction.
We encode our dialogue entries via a pretrained NLP Model. Previously, such models have already been used[18] for text anonymization tasks. Embeddings preserve the meaning of the text, which was shown in[34]. Pre-trained text embeddings can capture text sentiment and improve text-to-speech models[19].
6 Limitations
The study used data from only one company, which may limit the generalizability of the results to other organizations. In addition, the data was subject to de-identification, which limits the possibility of using models trained on this dataset outside of it. Also, the data analysis results can not be generalized. In other words, based on this dataset, it is impossible to draw conclusions regarding specific regions and market characteristics or perform deep text analytics. However, within the dataset, the data is consistent, which allows us to draw correct conclusions about the performance of multimodal or unimodal methods for working with sequences. It is also worth noting that the study was conducted on a sample of clients of a certain segment who had the opportunity to purchase certain products and does not cover all possible groups of consumers.
7 Ethics and Societal Implications
Our dataset is designed to facilitate the development of algorithms for analyzing complex event data, with the goal of enhancing the accuracy and efficiency of predictive models. Our event data does not contain characteristics such as gender, age, or race, which can be used to discriminate against certain groups of people. This makes the models trained on our dataset more ethical. However, it is important to note that companies must comply with legal grounds and obtain customer consent before collecting and analyzing additional data. It is also worth mentioning the need to ensure the security of data storage and transmission to prevent leakage of confidential information. Failure to comply with these conditions may result in privacy violations.
8 Conclusion and future work
In this paper, we introduced the first large-scale publicly available multimodal banking dataset, MBD333https://huggingface.co/datasets/ai-lab/MBD, which contains anonymized sequential data (bank transactions, geo position, and technical support dialogues) for 1M bank clients. It can be used to create industrial multimodal models for event sequences to predict client needs. Using this dataset, we present the novel benchmark based on MBD with two practically important tasks: campaigning (purchase prediction for the next month) and matching different modalities of clients. Our experimental results demonstrate that even a simple fusion of multiple data sources improves the overall quality compared to single-modal baselines (Table 2). Also, we show that anonymization does not significantly impact the algorithms’ performance, so our dataset is appropriate for choosing the best model that can be implemented in production scenarios. As a result, our work opens new horizons for developing practically relevant large-scale multimodal algorithms for event sequences.
In the future, expanding the range of tasks with our dataset is possible. For example, it is possible to estimate the multimodal techniques of the next-event prediction. Moreover, or baseline methods are based on late fusion, so it is very important to study more sophisticated models that consider possible mutual influence between modalities.
Appendix A Dataset Statistics
| Features | Number of categories | Description |
| amount | - | transaction amount |
| event time | - | transaction time |
| currency | 15 | transfer currency |
| event type | 56 | transaction type |
| event subtype | 62 | transaction subtype |
| src type11 | 101 | sender field type 1 |
| src type12 | 536 | sender field subtype 1 |
| dst type11 | 123 | receiver field type 1 |
| dst type12 | 649 | receiver field subtype 1 |
| src type21 | 40131 | sender field type 2 |
| src type22 | 87 | sender field subtype 2 |
| src type31 | 2389 | sender field type 3 |
| src type32 | 88 | sender field subtype 3 |
Appendix B Detailed experimental results
| methods | mean | target_1 | target_2 | target_3 | target_4 |
| DialogLast | 0.590 0.001 | 0.633 0.001 | 0.604 0.005 | 0.549 0.002 | 0.576 0.002 |
| DialogLast+GeoAggregation | 0.603 0.002 | 0.632 0.001 | 0.629 0.007 | 0.558 0.001 | 0.592 0.002 |
| DialogLast+GeoCoLES | 0.632 0.002 | 0.641 0.001 | 0.688 0.008 | 0.580 0.002 | 0.619 0.003 |
| DialogLast+GeoTabGPT | 0.648 0.002 | 0.654 0.001 | 0.715 0.008 | 0.591 0.003 | 0.632 0.003 |
| DialogLast+GeoTabBERT | 0.628 0.001 | 0.641 0.002 | 0.683 0.006 | 0.579 0.002 | 0.610 0.006 |
| DialogLast+TrxAggregation | 0.731 0.001 | 0.701 0.001 | 0.816 0.003 | 0.673 0.001 | 0.735 0.002 |
| DialogLast+TrxAggregation+GeoAggregation | 0.729 0.001 | 0.694 0.001 | 0.812 0.003 | 0.670 0.001 | 0.738 0.003 |
| DialogLast+TrxCoLES | 0.715 0.002 | 0.692 0.002 | 0.800 0.005 | 0.645 0.002 | 0.723 0.005 |
| DialogLast+TrxCoLES+GeoCoLES | 0.720 0.002 | 0.689 0.003 | 0.815 0.003 | 0.644 0.002 | 0.734 0.005 |
| DialogLast+TrxTabGPT | 0.739 0.001 | 0.683 0.001 | 0.820 0.005 | 0.685 0.001 | 0.767 0.005 |
| DialogLast+TrxTabGPT+GeoTabGPT | 0.749 0.002 | 0.690 0.001 | 0.838 0.005 | 0.687 0.001 | 0.780 0.005 |
| DialogLast+TrxTabBERT | 0.710 0.006 | 0.686 0.004 | 0.806 0.004 | 0.628 0.006 | 0.720 0.012 |
| DialogLast+TrxTabBERT+GeoTabBERT | 0.716 0.005 | 0.683 0.003 | 0.821 0.003 | 0.627 0.006 | 0.731 0.009 |
| DialogMean | 0.604 0.001 | 0.636 0.002 | 0.629 0.001 | 0.564 0.001 | 0.587 0.001 |
| DialogMean+GeoAggregation | 0.613 0.001 | 0.635 0.002 | 0.648 0.001 | 0.568 0.001 | 0.601 0.002 |
| DialogMean+GeoCoLES | 0.638 0.002 | 0.643 0.002 | 0.699 0.005 | 0.585 0.001 | 0.626 0.001 |
| DialogMean+GeoTabGPT | 0.653 0.001 | 0.656 0.001 | 0.725 0.005 | 0.595 0.002 | 0.638 0.003 |
| DialogMean+GeoTabBERT | 0.635 0.002 | 0.643 0.002 | 0.696 0.004 | 0.584 0.001 | 0.617 0.006 |
| DialogMean+TrxAggregation | 0.732 0.001 | 0.700 0.001 | 0.816 0.003 | 0.673 0.001 | 0.739 0.001 |
| DialogMean+TrxAggregation+GeoAggregation | 0.729 0.001 | 0.694 0.001 | 0.812 0.004 | 0.670 0.001 | 0.741 0.002 |
| DialogMean+TrxCoLES | 0.715 0.002 | 0.692 0.002 | 0.800 0.004 | 0.645 0.002 | 0.725 0.004 |
| DialogMean+TrxCoLES+GeoCoLES | 0.721 0.002 | 0.689 0.003 | 0.813 0.003 | 0.644 0.002 | 0.737 0.004 |
| DialogMean+TrxTabGPT | 0.739 0.001 | 0.682 0.001 | 0.819 0.005 | 0.684 0.001 | 0.769 0.004 |
| DialogMean+TrxTabGPT+GeoTabGPT | 0.748 0.001 | 0.689 0.001 | 0.837 0.004 | 0.686 0.001 | 0.782 0.004 |
| DialogMean+TrxTabBERT | 0.711 0.006 | 0.685 0.004 | 0.806 0.005 | 0.629 0.005 | 0.726 0.011 |
| DialogMean+TrxTabBERT+GeoTabBERT | 0.717 0.005 | 0.682 0.003 | 0.821 0.004 | 0.628 0.006 | 0.735 0.009 |
| GeoAggregation | 0.554 0.001 | 0.540 0.001 | 0.584 0.002 | 0.534 0.001 | 0.559 0.001 |
| GeoCoLES | 0.601 0.004 | 0.565 0.004 | 0.668 0.011 | 0.571 0.003 | 0.600 0.003 |
| GeoTabGPT | 0.622 0.001 | 0.589 0.001 | 0.700 0.008 | 0.586 0.003 | 0.615 0.004 |
| GeoTabBERT | 0.596 0.002 | 0.566 0.003 | 0.663 0.010 | 0.570 0.003 | 0.585 0.007 |
| TrxAggregation | 0.783 0.001 | 0.743 0.001 | 0.825 0.002 | 0.764 0.001 | 0.801 0.002 |
| TrxAggregation+GeoAggregation | 0.774 0.001 | 0.733 0.001 | 0.817 0.003 | 0.756 0.001 | 0.789 0.003 |
| TrxCoLES | 0.772 0.002 | 0.734 0.003 | 0.813 0.004 | 0.747 0.002 | 0.793 0.003 |
| TrxCoLES+GeoCoLES | 0.772 0.002 | 0.729 0.003 | 0.825 0.006 | 0.740 0.002 | 0.795 0.003 |
| TrxTabGPT | 0.796 0.000 | 0.746 0.001 | 0.837 0.004 | 0.778 0.001 | 0.825 0.004 |
| TrxTabGPT+GeoTabGPT | 0.798 0.001 | 0.743 0.001 | 0.850 0.003 | 0.772 0.001 | 0.827 0.004 |
| TrxTabBERT | 0.754 0.011 | 0.707 0.019 | 0.815 0.006 | 0.717 0.012 | 0.778 0.012 |
| TrxTabBERT+GeoTabBERT | 0.758 0.010 | 0.707 0.016 | 0.831 0.006 | 0.713 0.011 | 0.781 0.010 |
| methods | mean | target_1 | target_2 | target_3 | target_4 |
| DialogLast | 0.590 0.001 | 0.633 0.001 | 0.604 0.005 | 0.549 0.002 | 0.576 0.002 |
| DialogLast+GeoAggregation | 0.636 0.001 | 0.603 0.001 | 0.649 0.003 | 0.645 0.001 | 0.648 0.002 |
| DialogLast+GeoCoLES | 0.647 0.002 | 0.615 0.002 | 0.660 0.006 | 0.650 0.002 | 0.663 0.003 |
| DialogLast+GeoTabGPT | 0.654 0.002 | 0.631 0.001 | 0.673 0.005 | 0.652 0.002 | 0.662 0.002 |
| DialogLast+GeoTabBERT | 0.642 0.003 | 0.613 0.002 | 0.655 0.009 | 0.646 0.001 | 0.655 0.004 |
| DialogLast+TrxAggregation | 0.788 0.001 | 0.749 0.001 | 0.826 0.003 | 0.772 0.001 | 0.805 0.004 |
| DialogLast+TrxAggregation+GeoAggregation | 0.787 0.001 | 0.746 0.001 | 0.826 0.004 | 0.773 0.001 | 0.804 0.002 |
| DialogLast+TrxCoLES | 0.776 0.002 | 0.739 0.003 | 0.814 0.004 | 0.753 0.002 | 0.797 0.003 |
| DialogLast+TrxCoLES+GeoCoLES | 0.777 0.002 | 0.739 0.004 | 0.816 0.004 | 0.755 0.003 | 0.798 0.002 |
| DialogLast+TrxTabGPT | 0.805 0.001 | 0.775 0.001 | 0.842 0.002 | 0.778 0.002 | 0.823 0.005 |
| DialogLast+TrxTabGPT+GeoTabGPT | 0.802 0.001 | 0.764 0.001 | 0.845 0.002 | 0.777 0.001 | 0.820 0.004 |
| DialogLast+TrxTabBERT | 0.764 0.009 | 0.715 0.019 | 0.817 0.007 | 0.734 0.007 | 0.789 0.008 |
| DialogLast+TrxTabBERT+GeoTabBERT | 0.765 0.009 | 0.714 0.019 | 0.819 0.006 | 0.738 0.007 | 0.789 0.008 |
| DialogMean | 0.604 0.001 | 0.636 0.002 | 0.629 0.001 | 0.564 0.001 | 0.587 0.001 |
| DialogMean+GeoAggregation | 0.642 0.001 | 0.605 0.002 | 0.658 0.002 | 0.650 0.001 | 0.654 0.001 |
| DialogMean+GeoCoLES | 0.653 0.001 | 0.618 0.002 | 0.670 0.002 | 0.656 0.001 | 0.669 0.001 |
| DialogMean+GeoTabGPT | 0.661 0.001 | 0.633 0.001 | 0.681 0.004 | 0.657 0.002 | 0.671 0.002 |
| DialogMean+GeoTabBERT | 0.648 0.003 | 0.613 0.003 | 0.666 0.005 | 0.652 0.002 | 0.662 0.004 |
| DialogMean+TrxAggregation | 0.788 0.000 | 0.749 0.001 | 0.825 0.003 | 0.773 0.001 | 0.804 0.001 |
| DialogMean+TrxAggregation+GeoAggregation | 0.787 0.001 | 0.746 0.001 | 0.825 0.002 | 0.773 0.001 | 0.804 0.003 |
| DialogMean+TrxCoLES | 0.776 0.002 | 0.739 0.003 | 0.814 0.004 | 0.753 0.002 | 0.798 0.003 |
| DialogMean+TrxCoLES+GeoCoLES | 0.777 0.002 | 0.739 0.002 | 0.815 0.003 | 0.755 0.003 | 0.799 0.002 |
| DialogMean+TrxTabGPT | 0.805 0.001 | 0.775 0.001 | 0.843 0.001 | 0.778 0.002 | 0.824 0.004 |
| DialogMean+TrxTabGPT+GeoTabGPT | 0.802 0.001 | 0.764 0.001 | 0.845 0.002 | 0.777 0.001 | 0.821 0.003 |
| DialogMean+TrxTabBERT | 0.765 0.009 | 0.715 0.019 | 0.817 0.006 | 0.735 0.006 | 0.792 0.007 |
| DialogMean+TrxTabBERT+GeoTabBERT | 0.766 0.009 | 0.714 0.019 | 0.819 0.006 | 0.738 0.006 | 0.792 0.007 |
| GeoAggregation | 0.554 0.001 | 0.540 0.001 | 0.584 0.002 | 0.534 0.001 | 0.559 0.001 |
| GeoCoLES | 0.601 0.004 | 0.565 0.004 | 0.668 0.011 | 0.571 0.003 | 0.600 0.003 |
| GeoTabGPT | 0.622 0.001 | 0.589 0.001 | 0.700 0.008 | 0.586 0.003 | 0.615 0.004 |
| GeoTabBERT | 0.596 0.002 | 0.566 0.003 | 0.663 0.010 | 0.570 0.003 | 0.585 0.007 |
| TrxAggregation | 0.780 0.005 | 0.743 0.001 | 0.824 0.001 | 0.762 0.001 | 0.791 0.017 |
| TrxAggregation+GeoAggregation | 0.779 0.004 | 0.740 0.001 | 0.828 0.002 | 0.762 0.001 | 0.787 0.013 |
| TrxCoLES | 0.772 0.002 | 0.734 0.003 | 0.813 0.004 | 0.746 0.001 | 0.793 0.003 |
| TrxCoLES+GeoCoLES | 0.772 0.002 | 0.734 0.004 | 0.814 0.004 | 0.749 0.002 | 0.792 0.002 |
| TrxTabGPT | 0.796 0.000 | 0.745 0.001 | 0.837 0.004 | 0.777 0.001 | 0.824 0.005 |
| TrxTabGPT+GeoTabGPT | 0.796 0.001 | 0.751 0.002 | 0.843 0.003 | 0.774 0.001 | 0.816 0.003 |
| TrxTabBERT | 0.754 0.011 | 0.707 0.019 | 0.815 0.006 | 0.717 0.012 | 0.778 0.012 |
| TrxTabBERT+GeoTabBERT | 0.756 0.011 | 0.707 0.019 | 0.816 0.005 | 0.722 0.010 | 0.778 0.012 |
| methods | mean | target_1 | target_2 | target_3 | target_4 |
| DialogLast | 0.586 0.001 | 0.602 0.001 | 0.622 0.004 | 0.554 0.001 | 0.567 0.002 |
| DialogLast+GeoAggregation | 0.600 0.001 | 0.605 0.000 | 0.648 0.004 | 0.560 0.001 | 0.585 0.002 |
| DialogLast+GeoCoLES | 0.625 0.003 | 0.615 0.001 | 0.700 0.005 | 0.577 0.003 | 0.609 0.003 |
| DialogLast+GeoTabGPT | 0.642 0.002 | 0.629 0.001 | 0.725 0.007 | 0.589 0.002 | 0.623 0.002 |
| DialogLast+GeoTabBERT | 0.629 0.001 | 0.616 0.001 | 0.709 0.006 | 0.577 0.002 | 0.613 0.001 |
| DialogLast+TrxAggregation | 0.732 0.002 | 0.685 0.001 | 0.826 0.004 | 0.688 0.001 | 0.731 0.003 |
| DialogLast+TrxAggregation+GeoAggregation | 0.731 0.002 | 0.681 0.001 | 0.825 0.005 | 0.684 0.001 | 0.736 0.003 |
| DialogLast+TrxCoLES | 0.714 0.001 | 0.675 0.001 | 0.806 0.004 | 0.658 0.002 | 0.715 0.003 |
| DialogLast+TrxCoLES+GeoCoLES | 0.720 0.001 | 0.673 0.002 | 0.822 0.004 | 0.657 0.002 | 0.728 0.002 |
| DialogLast+TrxTabGPT | 0.743 0.002 | 0.677 0.001 | 0.834 0.004 | 0.697 0.001 | 0.766 0.003 |
| DialogLast+TrxTabGPT+GeoTabGPT | 0.753 0.001 | 0.684 0.001 | 0.848 0.004 | 0.700 0.001 | 0.779 0.003 |
| DialogLast+TrxTabBERT | 0.710 0.005 | 0.669 0.003 | 0.814 0.006 | 0.640 0.008 | 0.715 0.007 |
| DialogLast+TrxTabBERT+GeoTabBERT | 0.718 0.004 | 0.670 0.003 | 0.828 0.005 | 0.641 0.008 | 0.733 0.006 |
| DialogMean | 0.595 0.002 | 0.600 0.001 | 0.633 0.006 | 0.566 0.001 | 0.580 0.002 |
| DialogMean+GeoAggregation | 0.607 0.001 | 0.604 0.000 | 0.657 0.004 | 0.572 0.000 | 0.596 0.002 |
| DialogMean+GeoCoLES | 0.630 0.002 | 0.615 0.001 | 0.703 0.006 | 0.586 0.002 | 0.618 0.003 |
| DialogMean+GeoTabGPT | 0.645 0.002 | 0.629 0.001 | 0.727 0.010 | 0.596 0.001 | 0.630 0.001 |
| DialogMean+GeoTabBERT | 0.634 0.002 | 0.616 0.002 | 0.713 0.006 | 0.586 0.002 | 0.621 0.002 |
| DialogMean+TrxAggregation | 0.732 0.001 | 0.684 0.001 | 0.824 0.001 | 0.688 0.001 | 0.732 0.002 |
| DialogMean+TrxAggregation+GeoAggregation | 0.731 0.001 | 0.679 0.000 | 0.822 0.002 | 0.684 0.001 | 0.737 0.002 |
| DialogMean+TrxCoLES | 0.713 0.001 | 0.674 0.001 | 0.804 0.002 | 0.659 0.002 | 0.717 0.003 |
| DialogMean+TrxCoLES+GeoCoLES | 0.719 0.001 | 0.672 0.002 | 0.819 0.003 | 0.658 0.002 | 0.729 0.002 |
| DialogMean+TrxTabGPT | 0.742 0.001 | 0.675 0.001 | 0.829 0.004 | 0.697 0.001 | 0.766 0.003 |
| DialogMean+TrxTabGPT+GeoTabGPT | 0.751 0.001 | 0.682 0.001 | 0.845 0.004 | 0.700 0.001 | 0.779 0.004 |
| DialogMean+TrxTabBERT | 0.709 0.004 | 0.668 0.003 | 0.812 0.008 | 0.642 0.007 | 0.717 0.006 |
| DialogMean+TrxTabBERT+GeoTabBERT | 0.718 0.003 | 0.668 0.003 | 0.826 0.006 | 0.643 0.007 | 0.734 0.006 |
| GeoAggregation | 0.555 0.001 | 0.539 0.000 | 0.590 0.002 | 0.533 0.001 | 0.560 0.001 |
| GeoCoLES | 0.598 0.004 | 0.568 0.003 | 0.663 0.005 | 0.568 0.007 | 0.593 0.005 |
| GeoTabGPT | 0.621 0.003 | 0.589 0.002 | 0.696 0.010 | 0.586 0.002 | 0.614 0.002 |
| GeoTabBERT | 0.603 0.002 | 0.573 0.003 | 0.672 0.007 | 0.570 0.004 | 0.598 0.004 |
| TrxAggregation | 0.788 0.001 | 0.743 0.003 | 0.831 0.002 | 0.777 0.001 | 0.800 0.002 |
| TrxAggregation+GeoAggregation | 0.778 0.002 | 0.733 0.002 | 0.822 0.004 | 0.767 0.001 | 0.790 0.002 |
| TrxCoLES | 0.774 0.002 | 0.734 0.002 | 0.812 0.004 | 0.759 0.002 | 0.790 0.003 |
| TrxCoLES+GeoCoLES | 0.775 0.001 | 0.730 0.002 | 0.827 0.004 | 0.751 0.003 | 0.792 0.002 |
| TrxTabGPT | 0.802 0.001 | 0.751 0.001 | 0.844 0.002 | 0.788 0.002 | 0.826 0.003 |
| TrxTabGPT+GeoTabGPT | 0.804 0.001 | 0.748 0.001 | 0.854 0.002 | 0.784 0.001 | 0.829 0.003 |
| TrxTabBERT | 0.762 0.004 | 0.717 0.006 | 0.819 0.004 | 0.734 0.006 | 0.778 0.006 |
| TrxTabBERT+GeoTabBERT | 0.766 0.004 | 0.717 0.006 | 0.831 0.005 | 0.729 0.007 | 0.786 0.005 |
| methods | mean | target_1 | target_2 | target_3 | target_4 |
| DialogLast | 0.586 0.001 | 0.602 0.001 | 0.622 0.004 | 0.554 0.001 | 0.567 0.002 |
| DialogLast+GeoAggregation | 0.646 0.001 | 0.614 0.001 | 0.659 0.003 | 0.654 0.001 | 0.657 0.001 |
| DialogLast+GeoCoLES | 0.660 0.001 | 0.633 0.002 | 0.675 0.004 | 0.661 0.001 | 0.671 0.003 |
| DialogLast+GeoTabGPT | 0.668 0.001 | 0.645 0.001 | 0.690 0.004 | 0.662 0.001 | 0.674 0.002 |
| DialogLast+GeoTabBERT | 0.662 0.001 | 0.633 0.002 | 0.680 0.003 | 0.662 0.001 | 0.675 0.002 |
| DialogLast+TrxAggregation | 0.792 0.002 | 0.752 0.001 | 0.829 0.001 | 0.780 0.001 | 0.805 0.006 |
| DialogLast+TrxAggregation+GeoAggregation | 0.791 0.001 | 0.750 0.001 | 0.829 0.005 | 0.782 0.001 | 0.803 0.001 |
| DialogLast+TrxCoLES | 0.782 0.001 | 0.746 0.003 | 0.814 0.003 | 0.765 0.002 | 0.802 0.003 |
| DialogLast+TrxCoLES+GeoCoLES | 0.783 0.001 | 0.745 0.004 | 0.819 0.003 | 0.767 0.001 | 0.803 0.002 |
| DialogLast+TrxTabGPT | 0.810 0.001 | 0.779 0.001 | 0.846 0.003 | 0.789 0.002 | 0.827 0.004 |
| DialogLast+TrxTabGPT+GeoTabGPT | 0.808 0.001 | 0.770 0.001 | 0.849 0.003 | 0.790 0.001 | 0.824 0.004 |
| DialogLast+TrxTabBERT | 0.773 0.002 | 0.730 0.006 | 0.822 0.005 | 0.749 0.004 | 0.792 0.003 |
| DialogLast+TrxTabBERT+GeoTabBERT | 0.776 0.003 | 0.729 0.006 | 0.827 0.004 | 0.752 0.004 | 0.794 0.003 |
| DialogMean | 0.595 0.002 | 0.600 0.001 | 0.633 0.006 | 0.566 0.001 | 0.580 0.002 |
| DialogMean+GeoAggregation | 0.649 0.001 | 0.614 0.001 | 0.665 0.002 | 0.656 0.001 | 0.662 0.001 |
| DialogMean+GeoCoLES | 0.663 0.001 | 0.632 0.002 | 0.680 0.004 | 0.663 0.000 | 0.675 0.002 |
| DialogMean+GeoTabGPT | 0.670 0.001 | 0.645 0.000 | 0.694 0.004 | 0.664 0.001 | 0.678 0.001 |
| DialogMean+GeoTabBERT | 0.664 0.001 | 0.633 0.001 | 0.682 0.002 | 0.664 0.001 | 0.678 0.001 |
| DialogMean+TrxAggregation | 0.792 0.002 | 0.752 0.002 | 0.828 0.002 | 0.781 0.001 | 0.807 0.006 |
| DialogMean+TrxAggregation+GeoAggregation | 0.792 0.002 | 0.750 0.002 | 0.829 0.002 | 0.782 0.001 | 0.807 0.005 |
| DialogMean+TrxCoLES | 0.781 0.001 | 0.745 0.003 | 0.814 0.002 | 0.765 0.002 | 0.802 0.003 |
| DialogMean+TrxCoLES+GeoCoLES | 0.783 0.001 | 0.744 0.004 | 0.819 0.003 | 0.767 0.002 | 0.802 0.002 |
| DialogMean+TrxTabGPT | 0.810 0.002 | 0.779 0.001 | 0.847 0.004 | 0.789 0.001 | 0.828 0.003 |
| DialogMean+TrxTabGPT+GeoTabGPT | 0.808 0.001 | 0.770 0.001 | 0.848 0.003 | 0.790 0.001 | 0.826 0.003 |
| DialogMean+TrxTabBERT | 0.773 0.003 | 0.730 0.006 | 0.822 0.004 | 0.749 0.004 | 0.791 0.003 |
| DialogMean+TrxTabBERT+GeoTabBERT | 0.775 0.003 | 0.728 0.005 | 0.827 0.003 | 0.752 0.004 | 0.794 0.004 |
| GeoAggregation | 0.555 0.001 | 0.539 0.000 | 0.590 0.002 | 0.533 0.001 | 0.560 0.001 |
| GeoCoLES | 0.598 0.004 | 0.568 0.003 | 0.663 0.005 | 0.568 0.007 | 0.593 0.005 |
| GeoTabGPT | 0.621 0.003 | 0.589 0.002 | 0.696 0.010 | 0.586 0.002 | 0.614 0.002 |
| GeoTabBERT | 0.603 0.002 | 0.573 0.003 | 0.672 0.007 | 0.570 0.004 | 0.598 0.004 |
| TrxAggregation | 0.783 0.002 | 0.741 0.003 | 0.828 0.003 | 0.770 0.004 | 0.792 0.007 |
| TrxAggregation+GeoAggregation | 0.783 0.002 | 0.740 0.002 | 0.829 0.003 | 0.771 0.001 | 0.792 0.011 |
| TrxCoLES | 0.773 0.002 | 0.734 0.002 | 0.812 0.004 | 0.758 0.002 | 0.790 0.003 |
| TrxCoLES+GeoCoLES | 0.775 0.002 | 0.734 0.002 | 0.815 0.004 | 0.760 0.002 | 0.789 0.003 |
| TrxTabGPT | 0.802 0.001 | 0.751 0.001 | 0.844 0.002 | 0.787 0.001 | 0.825 0.003 |
| TrxTabGPT+GeoTabGPT | 0.800 0.001 | 0.752 0.001 | 0.846 0.005 | 0.785 0.002 | 0.817 0.006 |
| TrxTabBERT | 0.762 0.004 | 0.717 0.006 | 0.819 0.004 | 0.734 0.006 | 0.777 0.006 |
| TrxTabBERT+GeoTabBERT | 0.764 0.004 | 0.716 0.006 | 0.823 0.004 | 0.737 0.006 | 0.780 0.005 |
| Recall@1 | Recall@50 | Recall@100 | |
| Trx2Geo | 0.006 0.0003 | 0.1967 0.002 | 0.303 0.004 |
| Geo2Trx | 0.004 0.0003 | 0.1624 0.002 | 0.262 0.004 |
| Trx2Dial | 0.00004 0.00001 | 0.0001 0.00005 | 0.003 0.0001 |
| Dial2Trx | 0.00003 0.00001 | 0.001 0.00005 | 0.003 0.0001 |
| Dial2Geo | 0.00002 0.000001 | 0.0006 0.00001 | 0.001 0.0001 |
| Geo2Dial | 0.00002 0.000001 | 0.0005 0.00001 | 0.001 0.0001 |

### v2

@@
 Multimodal Banking Dataset: Understanding Client Needs through Event Sequences
-Abstract
-Financial organizations collect a huge amount of data about clients that typically has a temporal (sequential) structure and is collected from various sources (modalities).
-Due to privacy issues, there are no large-scale open-source multimodal datasets of event sequences, which significantly limits the research in this area.
-In this paper, we present the industrial-scale publicly available multimodal banking dataset, MBD, that contains more than 1.5M corporate clients with several modalities: 950M bank transactions, 1B geo position events, 5M embeddings of dialogues with technical support and monthly aggregated purchases of four bank’s products.
-All entries are properly anonymized from real proprietary bank data.
-Using this dataset, we introduce a novel benchmark with two business tasks: campaigning (purchase prediction in the next month) and matching of clients.
-We provide numerical results that demonstrate the superiority of our multi-modal baselines over single-modal techniques for each task.
-As a result, the proposed dataset can open new perspectives and facilitate the future development of practically important large-scale multimodal algorithms for event sequences.
+Abstract.
+Financial organizations collect a huge amount of temporal (sequential) data about clients, which is typically collected from multiple sources (modalities).
+Despite the urgent practical need, developing deep learning techniques suitable to handle such data is limited by the absence of large open-source multi-source real-world datasets of event sequences.
+To fill this gap, which is mainly caused by security reasons, we present the first industrial-scale publicly available multimodal banking dataset, MBD, that contains information on more than 2M corporate clients of a large bank.
+Clients are represented by several data sources: 950M bank transactions, 1B geo position events, 5M embeddings of dialogues with technical support, and monthly aggregated purchases of four bank products.
+All entries are properly anonymized from real proprietary bank data, and the experiments confirm that our anonymization still saves all significant information for introduced downstream tasks.
+Moreover, we introduce a novel multimodal benchmark suggesting several important practical tasks, such as future purchase prediction and modality matching.
+The benchmark incorporates our MBD and two public financial datasets.
+We provide numerical results for the state-of-the-art event sequence modeling techniques including large language models and demonstrate the superiority of fusion baselines over single-modal techniques for each task.
+Thus, MBD provides a valuable resource for future research in financial applications of multimodal event sequence analysis.
 HuggingFace Link: https://huggingface.co/datasets/ai-lab/MBD
 Github Link: https://github.com/Dzhambo/MBD
-1 Introduction
-The ever-growing digital footprint of consumers has led to an explosion of data collected by businesses across all industries.
-Financial institutions, in particular, gather a vast amount of information about their clients over time, encompassing various aspects of their financial activities.
-This data, spanning extended periods, is typically annotated with temporal information, forming what is known as event sequences [8, 14, 23, 37].
-Description of an event usually contains several heterogeneous fields, numerical and categorical, and processing of event sequence is a separate problem in machine learning different from techniques applied to time series or texts.
-An important property of event sequences is that these data are often gathered from multiple sources or channels, rendering multimodal.
-The comprehensive analysis of such data in a joint way is critical for informed decision-making processes within enterprises.
-The scarcity of publicly available datasets in this domain hampers research and development efforts.
-Indeed, the significant progress in multimodal data for natural language processing, computer vision, audio, etc., is due to the appearance of large-scale datasets, such as LAION, DataComp, etc.
-[16].
-Unfortunately, though several datasets of event sequences are used in research, e.g., credit card transactions [30] or MIMIC [21], they are either small or contain only one modality.
-Thus, a gap exists in methodologies designed to tackle the complexity of multimodal event sequence data.
-To bridge this gap, this paper introduces the Multimodal Banking Dataset (MBD), an unprecedented open-source resource encompassing extensive multichannel event sequence data of banking corporate clients.
-This dataset is the largest of its kind, featuring detailed records of approximately 1.5 million clients across four distinct modalities: money transfers (about 950 million events), geo position data (around 1 billion events), technical support dialog embeddings (approximately 5 million entries), and monthly aggregated bank product purchases categorized into four types.
-Each modality encompasses roughly one or two year of historical, time-annotated data, making it a rich resource for analyzing the dynamics of client behavior over time.
-The MBD dataset enables the exploration of several critical business problems in a multimodal context.
-Specifically, it facilitates research in such tasks as purchase prediction for the following month (campaigning), and matching of different modalities for the same clients.
-These problems can all be addressed by leveraging the temporal and spatial attributes inherent in the dataset.
-Moreover, we provide benchmarks for these tasks using the MBD.
-Importantly, all the data are anonymized to protect client privacy, and our research confirms that the anonymization process preserves the consistency of model performance between original and anonymized data.
-This paper underscores the potential of MBD as a foundational tool for advancing research in multimodal sequential data analysis, fostering innovation, and enhancing decision-making processes in the financial sector.
-2 Dataset
-Modern innovative banking institutions actively develop AI technologies for customizing their human-oriented technologies and making everyday decisions.
-A superior level of technologies will lead to new cases of customers’ experience, which should form a competitive advantage of services provided, including the speed, accuracy, and price of customer services, including personal credit conditions, individual finance strategy, etc.
+1. Introduction
+In the key tasks in the banking industry, such as campaigning, fraud detection, credit risk assessment, customer segmentation, and personalized recommendations, the performance heavily depends on the processing of the client’s financial and non-financial activities, e.g., transactions, dialogues with sales managers, product purchase history, etc.
+(Babaev et al., 2022; Rosset et al., 2001).
+This data, spanning extended periods, is typically annotated with temporal information, forming what is known as event sequences (Udovichenko et al., 2024; Yeshchenko and Mendling, 2022; Kolosnjaji et al., 2016; Weiss and Hirsh, 1998; Guo et al., 2020).
+An event is described by several heterogeneous fields, numerical and categorical.
+An essential property of event sequences is that these data are often gathered from multiple sources or channels, rendering them multimodal.
+The success of financial organizations strongly depends on their availability to deeply analyze such multi-source heterogeneous event sequences for building classification and predictive models.
+However, existing multimodal models (Xu et al., 2023; Zhang and Yan, 2023) cannot be directly applied to such event/tabular data due to their significant difference from audio, images, texts, and regular time series.
+Unfortunately, despite the urgent business needs, the progress in the development of multimodal techniques for multi-source event sequences is limited by the absence of large-scale datasets.
+Indeed, though several datasets of event sequences are used in research, e.g., credit card transactions (Padhi et al., 2021) or MIMIC (Johnson et al., 2023), they are either small or contain only one modality.
+Thus, tackling the complexity of multimodal event sequence data is still very challenging.
+To bridge this gap caused mainly by security reasons in the fintech industry, this paper introduces the Multimodal Banking Dataset (MBD), an unprecedented open-source resource encompassing extensive multichannel event sequence data of banking corporate clients.
+It is the largest dataset of its kind, featuring detailed records of approximately 2 million clients across four distinct modalities: money transfers (about 950 million events), geo position data (around 1 billion events), technical support dialog embeddings (approximately 5 million entries), and monthly aggregated bank product purchases categorized into four types.
+Each modality encompasses roughly one or two years of historical, time-annotated data, making it a rich resource for analyzing the dynamics of client behavior over time.
+The MBD dataset enables the research of several critical business problems in a multimodal context, such as future purchase prediction (campaigning) and matching different modalities for the same clients.
+To support these tasks, we provide benchmarks using the MBD and other existing financial datasets of much smaller size.
+To ensure client privacy, all data in MBD has been rigorously anonymized, allowing researchers to work with the dataset without compromising confidentiality.
+Moreover, our experiments confirm that our anonymization procedure preserves the consistency of model performance between original and anonymized data.
+2. Related Works
+2.1. Financial data
+Banks’ multitude of services and processes generates a variety of data that can be considered as modalities.
+Early works (Moro et al., 2014; Mancisidor et al., 2021) use feature processing techniques that remove multimodal complexity and present data in tabular form.
+A wide range of financial aggregates are presented in the American Express dataset (ame, [n.
+d.]) as a sequence of historical slices.
+The development of deep learning methods has led to the ability to work with complexly structured data, such as sequences of events.
+Quite a few datasets (age, [n.
+d.]; ros, [n.
+d.]; alp, [n.
+d.]), mostly unimodal, presented mainly at ML competitions.
+To work with such data, both supervised (Ala’raj et al., 2022; Babaev et al., 2019; Wang and Xiao, 2022) and unsupervised methods (Padhi et al., 2021; Bazarova et al., 2025; Skalski et al., 2023) are used.
+A multimodal financial sequential dataset was introduced in DataFusion 2022 competition (dat, [n.
+d.]).
+There are two sequential modalities, transaction and web clickstream, and two downstream tasks: matching and education level prediction.
+However, this is an extremely small dataset of 22K clients, and no accurate baseline model is available.
+2.2. Other event sequence domains
+Temporal point process (Mei and Eisner, 2017; Zhuzhel et al., 2023) model streams of discrete events in continuous time by constructing a neurally multivariate point process.
+The authors use a large collection of single-modal datasets from different types of modalities: Media (Retweets (Zhao et al., 2015), MemeTrack (Leskovec and Krevl, 2014), Amazon (ama, 2018), IPTV (Luo et al., 2014)), Medical (MIMIC-II (Johnson et al., 2016)), Social (Stack Overflow (Leskovec and Krevl, 2014), Linkedin (Xu and Zha, 2017)) and Financial (Transaction (Fursov et al., 2021)) data.
+Multimodal medical record datasets, MIMIC-IV (Johnson et al., 2023) is used in EventStreamGPT (McDermott et al., 2024).
+The authors propose a GPT-like approach for continuous-time event sequences.
+This dataset’s structure is close to our MDB.
+However, unlike medical information, financial data contains longer chains of events and more regular patterns, and individual transactions are less informative.
+2.3. Geostream and dialogues
+Geodata is used for various tasks.
+One of the uses of geo is a visualization of analytics on a map (Hao et al., 2011).
+However, geo is used here not as a separate modality but as additional tags to the mainstream of tweets.
+Mobile marketers (Baye et al., 2024) use geo-targeting for pricing and send personalized recommendations.
+In (Verma et al., 2020), geo hashes are used for user mobility detection and prediction.
+We encode our dialogue entries via a pretrained NLP model.
+Such models have already been used (Hassan et al., 2019) for text anonymization tasks.
+Embeddings preserve the meaning of the text, which was shown in (Vaswani et al., 2017).
+Pre-trained text embeddings can capture text sentiment and improve text-to-speech models (Hayashi et al., 2019).
+3. Proposed Dataset
+3.1. Overview of the MBD dataset
+Pipeline for processing original data sources for solving campaigning task.
+| Dataset | # Clients | Downstream Tasks | # Events | Class Balance | Modalities | ||||
+| Datafusion (dat, [n. d.]) | 22K |
+146M | Imbalanced | Transactions, Clickstream | ||||
+| Alphabattle (alp, [n. d.]) | 1.5M | Binary classification | 443M | Imbalanced | Transactions | ||||
+| Age (age, [n. d.]) | 50K | Multiclass classification | 44M | Balanced | Transactions | ||||
+| Rosbank (ros, [n. d.]) | 10K | Binary classification | 1M | Imbalanced | Transactions | ||||
+| Credit Card Transaction (Padhi et al., 2021) | 2K |
+2M | Highly Imbalanced | Transactions | ||||
+| MBD-small (ours) | 70K |
+80M | Imbalanced |
+| MBD (ours) | 2M |
+2B | Highly Imbalanced |
+Modern innovative banking institutions actively develop AI technologies to customize their human-oriented technologies and make everyday decisions.
+A superior level of technology will lead to new customer experience cases, which should form a competitive advantage for the services provided, including the speed, accuracy, and price of customer services, including personal credit conditions, individual finance strategy, etc.
 One of the main benefits of using AI is the ability to analyze large amounts of customer data.
@@
 In addition, the data is considered to have commercial value, and companies do not want to disclose it.
-To mitigate these risks, it is required to take all necessary steps to ensure data security and remove any identifying information.
+To mitigate these risks, all necessary steps must be taken to ensure data security and remove any identifying information.
 This allows information to be shared without violating client confidentiality, but this procedure requires significant effort from engineers, managers, and lawyers.
-Though there exist a small number of properly anonymized banking datasets, such as credit card transactions [30] or AlphaBattle [2], to the best of our knowledge, there are no publicly available large multimodal temporal datasets for banks.
-Thus, in this paper, we introduced the first large-scale multimodal banking dataset to support future research on multimodal techniques for event sequences.
-In particular, we selected several practically important tasks, such as campaigning, i.e., prediction if a client would purchase some of four rather popular products in the next month.
+Though there exist a small number of properly anonymized banking datasets, such as credit card transactions (Padhi et al., 2021) or AlphaBattle (alp, [n.
+d.]), to the best of our knowledge, there are no publicly available large multimodal temporal datasets for banks.
+Thus, this paper introduced the first large-scale multimodal banking dataset to support future research on multimodal techniques for event sequences.
+In particular, we select several practically essential tasks, such as campaigning (Rosset et al., 2001), i.e., prediction if a client would purchase some of four rather popular products in the next month.
 Each client is described by sequences of typical bank data: transactions, geo positions where the customer used the bank application, and dialogues with technical support.
-These data sources ideally highlight the main difficulties in developing multimodal models, namely, asynchronous events in different modalities, various intensities of events, rare/irregular events, and even the absence of some modalities for many clients.
-Based on our dataset, the researchers will be able to fully take into account cross-modal connections of sequences from multiple sources at the level of individual events.
-The dataset was collected using the following procedure.
-At first, we selected a complete sample of clients for two year (2021, 2022) to be able to cover all seasons.
-Among all customers who had the opportunity to purchase at least one of four products during 2022, we randomly selected 2,186,230 clients, among which 1M customers are labeled by monthly aggregated purchases of each of four products in each month.
-For these clients, we collected 947,899,612 bank transactions, 1,117,213,760 geo position events, and 5,080,781 dialogues with technical support.
-Our data raise typical practical challenges for training multimodal models.
-For example, many clients do not have all three modalities simultaneously because they can never make a transaction, call tech support, or leave their geo trace while running the bank application.
-Next, all the data were properly anonymized to guarantee the confidentiality and privacy of customer information.
-As a result, it becomes impossible to recover back to real clients from our anonymous data.
-We will show in the experimental study that such anonymization still let us extract valuable information about clients.
-To demonstrate the complete transformation of data from various sources, Fig.
-1 briefly shows example data for each modality, and the temporal structure of the data for a campaigning model is presented.
-Let us introduce further details for each modality in the following subsections.
-2.1 Modalities
+These data sources highlight the main difficulties in developing multimodal models: asynchronous events in different modalities, various intensities of events, rare/irregular events, and the absence of some modalities for many clients.
+Based on our dataset, the researchers can study cross-modal connections of sequences from multiple sources at the level of individual events.
+Let us discuss the details of the dataset collection procedure.
+At first, we select a complete sample of clients for two years (2021 and 2022) to cover all seasons.
+Among all customers who had the opportunity to purchase at least one of four products during 2022, we randomly chose 2,186,230 clients, among which 1M customers are labeled by monthly aggregated purchases of each of four products in each month.
+For these clients, we collect 947,899,612 financial operations, 1,117,213,760 geo position events, and 5,080,781 dialogues with technical support.
+Many clients do not have all three modalities simultaneously because they can never make a transaction, call tech support, or leave their geo trace while running the bank application.
+Next, all the data are properly anonymized to guarantee the confidentiality and privacy of customer information.
+As a result, it is impossible to recover real clients from our anonymous data.
+However, we will experimentally show that such anonymization still allows us to extract valuable client information.
+We present a comparative analysis of various event sequence datasets alongside the MBD dataset in Table 1.
+Our dataset is substantially more comprehensive, offering more modalities, events, clients, and downstream tasks than other datasets.
+MBD incorporates diverse modalities, including bank transactions, geo-locations, and technical support dialogues, providing a richer and more realistic basis for analysis.
+Additionally, it supports a broader range of downstream tasks, which are detailed in the following sections, allowing for more sophisticated and flexible modeling approaches.
+Given the substantial computational resources required for research with our dataset, we have curated a smaller and more balanced subset of MBD (“MBD-small”) to facilitate experimentation under resource-constrained conditions.
+We included all clients with at least one positive target value from the last month.
+Then, we also sampled clients with all zeros in their targets at a ratio of 1:10.
+The balanced subset contains 69474 clients in total.
+This subset retains the key characteristics of the full dataset while significantly reducing computational overhead.
+To demonstrate data transformation from various sources, Fig.
+1 briefly shows examples for each modality, and the temporal data structure for a campaigning model is presented.
+Let us briefly describe each modality in the following subsection.
+Further details are provided in Appendix A, including the sequence length of event sequences and data samples.
+3.2. Modalities
 1.
-Bank transactional data are financial operations (events) carried out between different clients.
-The sequence of transactions can uniquely characterize the client [8], so this modality plays one of the most important roles in financial planning and recommendations.
-Thus, the main component of our MBD dataset is each client’s transactional history, represented by an event with a timestamp and various attributes of the anonymized counterparty.
+Bank transactional data are financial operations carried out between different clients.
+The main component of our MBD dataset is the client’s transactional history, represented by an event with a timestamp and various attributes of the anonymized counterparty.
 Clients have 638 transactions on average.
-Fig.
-2 in Appendix A contains the detailed distribution of clients for different lengths of the transaction history.
-The list of attributes is shown in Table 4.
+Collected over two years (2021 and 2022), the sequence of financial operations can uniquely characterize the client (Babaev et al., 2022), so this data source plays one of the most significant roles in planning and recommendations.
 2.
-Dialogues.
-The dialogue data are transcriptions of customer calls to technical support and negotiations between the client and its manager.
-It is an extremely important source of information about client needs and problems [11].
-The audio utterance is fed into a commercial Speech-to-Text algorithm.
+Dialog data consists of transcriptions from customer calls to technical support and negotiations between clients and their managers.
+We incorporate dialogues from key communication channels, including sales and service calls, which account for most interactions with bank customers.
+It is an extremely important source of information about client needs and problems (Bauman et al., 2024).
+The audio utterance is fed into a commercial Speech-to-Text library.
 Personal information, e.g., the client’s name, is detected in the text and masked.
 To further anonymize the dialogue, we feed its text into a pre-trained NLP model111 https://huggingface.co/ai-forever/ruBert-base and save the resulting embeddings of size 768 in dialogue modality.
-An example of dialogue data is presented in Fig.
-3 in Appendix A.
-Only 46% of customers contact support and have records of conversations, 98% of them have no more than 10 dialogues per year.
-Fig.
-4 contains the detailed distribution of clients for various numbers of dialogues.
+Only 46% of customers contact support and have records of conversations, and 98% of them have no more than 10 dialogues.
 3.
-Geostream data contains a sequence of geo-coordinates of a client.
+Geostream data contains a sequence of geo-coordinates of a client obtained throughout 2022.
 To anonymize this modality, the coordinates are encoded using geohashes222https://pypi.org/project/pygeohash/, a geocoding system that converts a geographic location into a short string.
@@
 It is possible to adjust the accuracy and size by removing characters from the end of the code.
-In our dataset, the coordinates were encoded with a precision of 4, 5, and 6 characters, representing cells of different sizes on the map.
-In our dataset, there are 43,999 distinct values of geohash_4, 347,698 numbers of geohash_5, and 2,264,404 most precise locations (geohash_6).
-An example of geo data is presented in Fig.
-5 in Appendix A.
-The clients have an average of 910 geo-points in the time period considered.
-Detailed information about the distribution of their length over the clients is available in Fig.
-6.
+In our dataset, the coordinates are encoded with a precision of 4, 5, and 6 characters, representing cells of different sizes on the map.
+As a result, there are 43,999 distinct values of geohash_4, 347,698 numbers of geohash_5, and 2,264,404 most precise locations (geohash_6).
 4.
 Products purchases.
-Our dataset is a valuable resource for analyzing the bank customers’ needs.
-In particular, it contains information about purchases of four different banking products for each month of 2022.
+Our dataset serves as a valuable resource for analyzing bank customers’ needs and optimizing the campaigning process.
+This critical task directly impacts the bank’s product volume and overall profitability.
+High-quality recommendations play a key role in enhancing the customer experience, making campaigning essential for business outcomes and customer satisfaction.
+Specifically, MBD includes monthly data on the purchases of four distinct banking products throughout 2022, providing a broader temporal scope that captures patterns beyond the pandemic’s peak.
+We concentrate on the most popular of these products, as internal analysis across various tasks using proprietary data consistently showed that these products provide a robust foundation for model selection.
+The insights from this data enable the development of models that demonstrate superior performance across a broad spectrum of related tasks.
 To predict a purchase in a certain month, it is necessary to take events (transactions, geo, dialogues) strictly before the beginning of this month.
-Therefore, the date range for the purchases dataset is shifted by 1 month, i.e., information is available from February 1, 2022, to January 31, 2023.
-The campaigning task is a multi-label classification problem, i.e., we store a binary label for each product that indicates whether it was purchased by a customer in a certain month.
-Example of data about purchases is shown in Fig.
-8 (Appendix A).
+Therefore, the purchase dataset’s date range is shifted by 1 month, i.e., information is available from February 1, 2022, to January 31, 2023.
+The campaigning task is a multi-label classification problem, i.e., we store a binary label for each product that indicates whether a customer purchases it in a certain month.
 The peculiarity of this dataset is its imbalance, which is specific to this type of business task: 81% of clients have no purchases, 15% have one, and the remaining 4% have two or more purchases.
 A historical overview over 12 months allows us to model the customer behavior dynamic and predict the date of purchase more accurately.
-2.2 Data anonymization
+3.3. Data anonymization
 Our dataset contains no personal or confidential information whatsoever.
 Nevertheless, the event sequences are detailed enough that it could be possible to compare individuals from the publicly accessible portion of the dataset with the original proprietary data.
-To mitigate this risk, noise has been introduced to the data, ensuring that such comparisons and identification are impossible.
-The noise patterns were selected by our bank’s internal security department.
+To mitigate this risk, noise is introduced to the data, ensuring that such comparisons and identification are impossible.
+Our bank’s internal security department selected the noise patterns.
 These patterns are applied locally, preserving the overall structure of the data.
-The specific parameters of the noise are not disclosed to prevent potential attacks on the dataset.
-All ID fields were hashed with a random salt.
-All categorical field values was mapped to enumerated indexes.
-Random noise was added to numerical fields.
-The noise was added to date fields with preserving hour of original date, which may be the cause the shuffle of local sequence.
-The dialogue embedding space was divided into regions, which were then shuffled.
-3 Benchmark
-In this paper, we formulate two practically important tasks using our new dataset: campaigning and client matching. This section contains details about baseline methods and testing protocols.
-3.1 Methods
-To establish performance baselines, we implement several widely-adopted architectures.
-Our approach prioritizes unsupervised and semi-supervised methods [10], enabling the training of a general-purpose encoder on unlabeled sequential data.
-Additionally, we incorporate supervised methods that allow for the immediate training of the encoder in a fully supervised manner.
-The following techniques are implemented in our benchmark to extract features for Transactions and Geostream modalities:
+The specific noise parameters are not disclosed to prevent potential attacks on the dataset.
+All ID fields are hashed with a random salt.
+All categorical field values are mapped to enumerated indexes.
+Random noise is added to numerical fields and dates, preserving the hour of the original date, which may cause the local sequence’s shuffle.
+The dialogue embedding space is divided into regions, which are then shuffled.
+In the experimental study for downstream tasks, we demonstrate that the applied anonymization techniques have a minimal impact on the ranking of models trained on our dataset.
+This indicates that the anonymization process preserves fundamental structural properties while retaining information essential for the downstream task, thereby ensuring the reliability of model evaluation.
+4. Benchmark
+In this paper, we introduce a benchmark for widely used event sequence datasets, incorporating practically important downstream tasks.
+This section provides a detailed description of the downstream tasks, baseline methods, and evaluation protocols.
+4.1. Datasets and downstream tasks
+We implement an out-of-fold validation protocol for each downstream task in every dataset.
+The client dataset is partitioned into five folds, with four folds used for training and the remaining fold reserved for testing.
+The training and testing sets are publicly available alongside the dataset, allowing future researchers to compare performance metrics.
+As each dataset in our benchmark exhibits label imbalance, ROC AUC is the most robust and informative evaluation metric due to its resilience to class imbalances.
+In real-world business, campaign effectiveness is measured by revenue, but conducting A/B tests for every ML model is impractical.
+Instead, ROC AUC is a reliable proxy metric for model comparison, with only the top performers advancing to A/B testing.
+This method was validated through real-world A/B tests and is recognized as the core evaluation metric by leading institutions, including one of the largest global banks.
+The choice of AUC for campaigning is supported by the reason that ranking models by their ROC curves are similar to comparing their non-response ratio at all possible cutoff points simultaneously (Liu et al., 2012; Rosset et al., 2001).
+Let us discuss the details of downstream tasks for each dataset in our benchmark.
 1.
-Aggregation Baseline that contains hand-crafted aggregation statistics [8]: 3,000 and 4 features in total for sequences of transactions and geodata, respectively.
-2.
-CoLES (Contrastive Learning for Event Sequences), a self-supervised contrastive model [8] specially developed to obtain representations of such event sequences as bank transactions.
-The sequence encoder is a GRU (Gated Recurrent Unit) with a hidden size 256.
-3.
-Two Tabular Transformers from IBM [30].
-The first model, TabBERT, adapts BERT to event sequences such as bank transactions.
-The second model, TabGPT, was initially proposed to generate synthetic tabular sequences.
-Both models extract 256-dimensional embeddings of an input event sequence.
-After, we pool output embeddings of the client in result embedding of size 1024, calculating min, max, mean, and std.
-To obtain representation of a sequence of dialogues, we implemented several straightforward techniques to aggregate the sequence of embeddings of each dialog of a client: 1) mean pooling of all embeddings; and 2) use only the most recent embedding for the date of interest.
-To demonstrate the potential of multimodal processing of event sequences, we implemented traditional fusion techniques, namely, 1) blending, i.e., a weighted sum of scores (estimates of class posterior probabilities) from single-modal classifier, and 2) late fusion, i.e., learning a classifier for concatenated embeddings from all modalities [20]
-For supervised methods, we use the following architectures: for unimodal data, we utilize an GRU encoder, while for multimodal data, we employ different GRU encoders for each modality and embeddings from all encoders are concatenated.
-The resulting neural network is trained using binary cross-entropy (BCE) loss in multilabel setting.
-For the GRU encoders we used hidden size 32.
-3.2 Tasks
-3.2.1 Campaigning task
+MBD.
+For our dataset, we introduce a campaigning downstream task (Pavlova, 2024; Balyemah et al., 2024; Hendriksen et al., 2020).
 In this task, it is required to predict the customer’s propensity to purchase four different products in the next month (Fig.
 1) given sequences of transactions, geo locations, and dialogues from the beginning of this month.
-Solutions to this problem are used to plan marketing campaigns and prepare sales communications through various communication channels with the client.
-We implement an out-of-fold validation protocol to conduct our experiments.
-The entire client dataset is divided into five folds: four folds are used for training, while the remaining fold is designated for testing.
-Both the training and testing sets are made publicly available alongside our dataset to enable future researchers to compare performance metrics.
-The baseline methods outlined earlier are applied as follows.
-First, we train our models using the training set.
-Considering the temporal structure of our target, we compute the embedding of each client’s history up to one month, focusing on the presence of the target product (Fig.
+Solutions to this problem are used to plan marketing campaigns and prepare sales communications through various channels with the client.
+The baseline methods outlined in the following section are applied.
+The multi-label classification models are trained using our training set.
+Considering the temporal structure of our target, we compute the embedding of each client’s history for up to one month, focusing on the presence of the target product (Fig.
 1).
-We then evaluate the model using the multi-label classification metric ROC AUC across the 12 months of 2022 and for four binary product labels.
-3.2.2 Multimodal Matching
-Multimodal matching [41] involves aligning and comparing different modalities to identify meaningful relationships or connections between them.
-Frequently, data from multiple sources for the same client are matched using predefined rules or heuristics, which may not always yield optimal results.
-To enhance the accuracy of this process, specialized identification algorithms are required to more precisely compare different modalities.
-For the matching task, we employed a framework analogous to CLIP [31].
-We utilized GRU encoders to embed pairs of samples from two input modalities, labeling them as either positive matches (i.e., data from the same client) or negative matches (i.e., data from different clients).
-The model was trained using the InfoNCE loss function [13], which maximizes similarity for positive pairs while minimizing it for negative pairs.
-To assess the model’s performance, we used Recall@1, Recall@50 and Recall@100 metrics.
-4 Experiments
+We then evaluate the model using the mean ROC AUC across four binary product labels over the 12 months of 2022.
+2.
+Datafusion.
+In this dataset, the downstream task is to predict the higher education attainment of bank clients.
+It involves analyzing two modalities (transaction histories and clickstream data) to accurately infer client’s educational background.
+The task is formulated as a binary classification problem, with 75% of the labels corresponding to clients with higher education.
+3.
+Alphabattle.
+We incorporate the large unimodal Alphabattle dataset alongside the multimodal datasets into our benchmark.
+This inclusion of data from various financial institutions aims to support more robust and reliable conclusions.
+In this dataset, the downstream task estimates the probability of a customer defaulting based on their historical card transaction behavior.
+This task is framed as a binary classification problem, with 2.7% of the labels representing clients who have defaulted.
+Although the dataset is unimodal, the downstream task remains highly relevant for financial institutions, offering critical insights into credit risk assessment in improving decision-making processes related to customer management and financial strategies.
+Multimodal matching.
+For the multimodal datasets, MBD and Datafusion, we propose a downstream task of multimodal matching (Zong et al., 2023; Huang et al., 2021; Taghibakhshi et al., 2023; Phan et al., 2017).
+It involves aligning and comparing modalities to identify meaningful relationships or connections.
+Data from multiple sources for the same client are frequently matched using predefined rules or heuristics, which may not always yield optimal results.
+To improve its accuracy, specialized identification algorithms are required to compare modalities more precisely.
+For the matching task, we employ a framework analogous to CLIP (Radford et al., 2021).
+We utilize GRU encoders to embed pairs of samples from two input modalities, labeling them as either positive (i.e., data from the same client) or negative matches (i.e., data from different clients).
+The model is trained using the InfoNCE loss function (Chen et al., 2020), which maximizes similarity for positive pairs while minimizing it for negative pairs.
+We use Recall@1, Recall@50, and Recall@100 metrics to assess the model’s performance.
+4.2. Methods
+To establish performance baselines, we implement several widely adopted architectures.
+Our approach prioritizes unsupervised and semi-supervised methods (Balestriero et al., 2023), enabling the training of a general-purpose encoder on unlabeled sequential data.
+Additionally, we incorporate supervised methods that allow for the immediate training of the encoder in a fully supervised manner.
+4.2.1. Unimodal approaches
+The following techniques are implemented in our benchmark to extract features from data.
+First, Aggregation Baseline that contains hand-crafted aggregation statistics (Babaev et al., 2022).
+Events are represented either numerically, such as transaction amount, or categorically, like event types.
+For numerical attributes, we apply aggregation functions (e.g., sum, mean, std) across all events in a sequence.
+Categorical attributes are grouped by unique values, aggregating numerical attributes using functions like count or mean.
+Second, CoLES (Contrastive Learning for Event Sequences), a self-supervised contrastive model (Babaev et al., 2022) specially developed to obtain representations of such event sequences as bank transactions.
+The sequence encoder is a GRU (Gated Recurrent Unit) with a hidden size 256.
+Despite its simplicity, CoLES continues to demonstrate top-level performance in event sequence analysis (Osin et al., 2024).
+Finally, two Tabular Transformers from IBM (Padhi et al., 2021): 1) TabBERT that adapts BERT to event sequences such as bank transactions, and 2) TabGPT that was initially proposed to generate synthetic tabular sequences.
+Both models extract 256-dimensional embeddings of an input event sequence.
+After that, we pool output embeddings of the client in result embedding of size 1024, calculating min, max, mean, and std.
+To obtain a representation of a sequence of dialogues, we borrow conventional techniques to aggregate the sequence of embeddings of each client’s dialog: 1) mean pooling of all embeddings, and 2) use only the most recent embedding for the date of interest.
+For unimodal supervised methods (Supervised RNN), we utilize GRU architectures (Babaev et al., 2019) with a hidden size of 32.
+The models are trained in a multi-label setting using binary cross-entropy (BCE) loss, ensuring effective optimization for tasks with multiple targets.
+4.2.2. Multimodal approaches
+To explore the potential of multimodal processing for event sequence analysis, we compare several fusion techniques: Blending, Late, and Early Fusion.
+Blending computes a weighted sum of class posterior probabilities from individual single-modal classifiers, effectively combining the predictions from each modality.
+In Late Fusion, embeddings from all data sources are concatenated and fed into a classifier.
+This technique allows the model to learn interactions between modalities after they have been individually processed (Huang et al., 2020).
+In supervised Late Fusion, we utilize separate GRU encoders for each modality, concatenating their embeddings to form a unified representation (Supervised RNN).
+Early Fusion combines representations from multiple modalities at the initial stages of the model, enabling joint processing of multimodal data.
+We employ the CrossTransformer approach (Zhang and Yan, 2023), which utilizes a cross-attention mechanism to integrate information across modalities efficiently.
+In our experiments, this method is applied within a supervised learning framework.
+4.2.3. LLM approaches
+In addition to the methods described above, we also explore the use of Large Language Models (LLMs) for multimodal data analysis.
+Leveraging their ability to process sequential data through universal text-based representations, LLMs offer a versatile approach to analyzing events across different modalities.
+We used the X-separated serialization (Fang et al., 2024) to represent transactions in text format.
+X-separated serialization formats the data such that each transaction is represented as a separate line (row), with individual features separated by tabs (columns).
+In an unimodal setup, we limited transactions to 256 events and four features (event_time, event_type, amount, src_type3).
+In the multimodal setup, we used 128 transactions and 64 geo events.
+Serialized transactions and geo hashes are placed one after the other and separated.
+We used a special delimiter to separate modalities.
+We applied LLMs on serialized data, considering each user as a distinct document.
+Transactional sequence embedding is obtained from the last hidden layer’s last token.
+No extra model pretraining or prompting has been applied.
+These embeddings were used to solve downstream tasks for each target.
+We also used PCA on the embeddings to reduce their dimensionality to 256, matching the baseline embeddings.
+Our findings indicate that applying PCA to text embeddings reduces dimensionality and improves model predictive performance on downstream tasks.
+We used a popular AutoML library, LightAutoML (Vakhrushev et al., 2021) to solve downstream tasks.
+It provides a fully automatic pipeline for model selection and hyperparameter tuning.
+| Model | MBD | Datafusion | Alphabattle | ||
+| Transactions | Geostream | Transactions | Clickstream | Transactions | |
+| Aggregation | 0.783 0.002 | 0.595 0.002 | 0.793 0.013 | 0.537 0.018 | 0.785 0.0010 |
+| CoLES | 0.773 0.002 | 0.598 0.004 | 0.784 0.012 | 0.641 0.013 | 0.793 0.0005 |
+| TabBERT | 0.762 0.004 | 0.603 0.002 | 0.762 0.014 | 0.590 0.026 | 0.778 0.0003 |
+| TabGPT | 0.802 0.002 | 0.621 0.003 | 0.766 0.013 | 0.618 0.016 | 0.775 0.0010 |
+| Supervised RNN | 0.819 0.002 | 0.540 0.012 | 0.712 0.016 | 0.563 0.011 | 0.792 0.0030 |
+| Dataset | Modalities | CoLES | TabGPT | TabBERT | Supervised RNN |
+| MBD | Trx | 0.773 0.002 | 0.802 0.001 | 0.762 0.004 | 0.819 0.002 |
+| Trx + Geo | 0.775 0.002 | 0.800 0.001 | 0.764 0.004 | 0.819 0.001 | |
+| Trx + Dialog | 0.781 0.002 | 0.810 0.002 | 0.773 0.003 | 0.821 0.0006 | |
+| Trx + Dialog + Geo | 0.783 0.002 | 0.808 0.001 | 0.775 0.003 | 0.824 0.001 | |
+| Private dataset | Trx | 0.772 0.002 | 0.796 0.000 | 0.754 0.011 | 0.813 0.002 |
+| Trx + Geo | 0.772 0.002 | 0.796 0.001 | 0.756 0.011 | 0.815 0.001 | |
+| Trx + Dialog | 0.776 0.002 | 0.805 0.001 | 0.765 0.009 | 0.817 0.0006 | |
+| Trx + Dialog + Geo | 0.777 0.002 | 0.802 0.001 | 0.766 0.009 | 0.821 0.001 | |
+| Datafusion | Trx | 0.784 0.012 | 0.766 0.013 | 0.762 0.014 | 0.712 0.015 |
+| Trx + Click | 0.785 0.011 | 0.766 0.011 | 0.761 0.012 | 0.703 0.008 |
+5. Experiments
 Our models, experiments, and training procedures were implemented in Python, leveraging PyTorch and PyTorch Lightning for deep learning tasks, and PySpark for distributed data processing.
-We trained the neural networks using NVIDIA V100 GPU, while the boosting models were trained on computational clusters equipped with 600 cores.
-The reported experiments, including extensive hyperparameters optimization, required approximately 500 hours of computation time.
-4.1 Model and training hyperparameters
-We employed the unsupervised baseline methods (CoLES, TabGPT, TabBERT, Aggregation) with default hyperparameters from the pytorch-lifestream framework.
-The PyTorch implementation of the Adam optimizer was utilized, with an initial learning rate of 0.001, coupled with the StepLR scheduler.
-Models were saved based on the lowest validation loss or the highest validation unsupervised metrics [33], evaluated after each training epoch, with training of 15 epochs.
-We employed 24-dimensional embeddings for categorical features and clipped the number of categories for features with many unique values.
-We applied either an identical mapping or a logarithmic transformation for numerical features.
-We used the gradient boosting algorithm available in PySpark ML for downstream tasks.
-4.2 Campaigning task
-In this Subsection, we provide experimental results for the campaigning task in unimodal and multimodal baselines.
+We trained the models on NVIDIA V100 GPU, while the gradient boosting models were trained on computational clusters equipped with 600 CPU cores.
+5.1. Model and training hyperparameters
+We employ the unsupervised baseline methods (CoLES, TabGPT, TabBERT, Aggregation) with default hyperparameters from the pytorch-lifestream framework.
+The Adam optimizer is utilized, with an initial learning rate of 0.001, coupled with the StepLR scheduler.
+Models are saved based on the lowest validation loss or the highest validation unsupervised metrics (Tsitsulin et al., 2023), evaluated after each training epoch, with training of 15 epochs.
+We employ 24-dimensional embeddings for categorical features and clipped the number of categories for features with many unique values.
+We apply either an identical mapping or a logarithmic transformation for numerical features.
+We use the gradient boosting algorithm from PySpark ML for downstream tasks.
+5.2. Downstream tasks
+5.2.1. MBD: campaigning task
+This Subsection contains experimental results for the campaigning task in the MBD dataset in unimodal and multimodal baselines.
 One of the main objectives of our paper is to provide a real data benchmark to facilitate the development of multimodal algorithms.
-To achieve this, we must demonstrate that an algorithm outperforming another on our public benchmark will similarly outperform it on real data.
-Table 1 shows that the transaction modality plays the most crucial role in achieving accurate classification.
-In contrast, dialogues and geostream, when used in a unimodal setting, performed only slightly better than a random estimator.
-However, as shown in Table 2, the predictive performance improved significantly in the multimodal setting through the integration of additional modalities.
+It is necessary to demonstrate that an algorithm outperforming another on our public benchmark will similarly outperform it on real data.
+Table 2 shows that the transaction modality is crucial in achieving accurate classification.
+In contrast, dialogues and geostream perform only slightly better than a random estimator when used in unimodal setting.
+However, as shown in Table 3, the predictive performance improved significantly in the multimodal setting by integrating additional modalities.
 The overall trend indicates a consistent improvement in validation metrics as more modalities are incorporated.
-Specifically, the multimodal late fusion approach enhanced predictive accuracy by 1–2% when adding a single data source to the transaction stream, and by up to 3% when utilizing all available data sources.
-Models trained on public datasets generally exhibit higher performance scores, which is an anticipated outcome of the data anonymization process.
-Nevertheless, the relative ranking of different methods remains consistent across both MBD and private datasets.
+Specifically, the multimodal late fusion approach enhances predictive accuracy by 1-1,5% when adding other data sources to the transaction stream.
+Model performance comparison on private and public data.
 Fig.
-7 in Appendix B demonstrates a strong correlation in performance metrics between the public and private datasets, with a Kendall-tau correlation coefficient of 0.94.
-Comprehensive results for each modality and all possible fusion combinations of multiple modalities are provided in Appendix B.
-We denote our modalities: geostream, transactions, and dialogues are represented as Geo, Trx, and Dialogs, respectively.
-We specify the method with the modality to indicate that a particular method was applied.
-For example, if embeddings from the CoLES model are used on transactions, we denote this as TrxCoLES.
-| Modalities | Methods | MBD | Private dataset |
-| Dialog | Mean | 0.595 0.002 | 0.604 0.001 |
-| Last | 0.586 0.001 | 0.590 0.001 | |
-| Supervised | 0.540 0.012 | 0.538 0.010 | |
-| Geo | Aggregation | 0.555 0.001 | 0.554 0.001 |
-| TabGPT | 0.621 0.003 | 0.622 0.001 | |
-| TabBERT | 0.603 0.002 | 0.596 0.002 | |
-| CoLES | 0.598 0.004 | 0.601 0.004 | |
-| Supervised | 0.566 0.002 | 0.563 0.002 | |
-| Trx | Aggregation | 0.783 0.002 | 0.791 0.017 |
-| TabGPT | 0.802 0.002 | 0.796 0.000 | |
-| TabBERT | 0.762 0.004 | 0.754 0.011 | |
-| CoLES | 0.773 0.002 | 0.772 0.002 | |
-| Supervised | 0.819 0.002 | 0.813 0.002 |
-| Methods | Modalities | MBD | Private dataset |
-| CoLES | Trx | 0.773 0.002 | 0.772 0.002 |
-| Trx + Geo | 0.775 0.002 | 0.772 0.002 | |
-| Trx + Dialog | 0.781 0.001 | 0.776 0.002 | |
-| Trx + Dialog + Geo | 0.783 0.001 | 0.777 0.002 | |
-| TabGPT | Trx | 0.802 0.001 | 0.796 0.000 |
-| Trx + Geo | 0.800 0.001 | 0.796 0.001 | |
-| Trx + Dialog | 0.810 0.002 | 0.805 0.001 | |
-| Trx + Dialog + Geo | 0.808 0.001 | 0.802 0.001 | |
-| Supervised | Trx | 0.819 0.002 | 0.813 0.002 |
-| Trx + Geo | 0.819 0.001 | 0.815 0.001 | |
-| Trx + Dialog | 0.821 0.0006 | 0.817 0.0006 | |
-| Trx + Dialog + Geo | 0.824 0.001 | 0.821 0.001 | |
-| TabBERT | Trx | 0.762 0.004 | 0.754 0.011 |
-| Trx + Geo | 0.764 0.004 | 0.756 0.011 | |
-| Trx + Dialog | 0.773 0.003 | 0.765 0.009 | |
-| Trx + Dialog + Geo | 0.775 0.003 | 0.766 0.009 |
-4.3 Multimodal matching task
-In this subsection, we present the results of our proposed multimodal matching benchmark, summarized in Table 3, which includes both transactions and dialogues.
+2 highlights a strong correlation between performance metrics on public and private datasets, with a Kendall-tau correlation coefficient of 0.94.
+This figure presents performance results derived from the extended experiments detailed in Appendix B (Tables 11, 13) for the campaigning task on both datasets.
+The private dataset is structurally equivalent to the public dataset, maintaining identical client composition, temporal coverage, and data volume, with the sole distinction being the application of anonymization.
+Hence, the anonymization process has minimal impact on model performance for the downstream task of campaigning.
+The consistency in relative ranking across both datasets (Table 3) underscores the reliability of our benchmark for advancing research in multimodal event sequence analysis.
+More detailed comparison of results on MBD and the private dataset and comprehensive results for each modality and all possible fusion combinations of multiple modalities are shown in Appendix B (Tables 10 - 13).
+Our modalities, namely, geostream, transactions, and dialogues, are denoted as Geo, Trx, and Dialogs.
+We use a clear notation to specify a method applied to a modality.
+For instance, if embeddings from the CoLES model are applied to transactions, we denote this as TrxCoLES.
+5.2.2. Datafusion: Higher Education
+In this subsection, we present the results of unimodal and multimodal experiments on the DataFusion dataset, as shown in Tables 2 and 3.
+While the dataset is small, leading to an insignificant increase in quality metrics when incorporating additional sources in multimodal settings, the results in both unimodal and multimodal configurations remain valuable as they contribute to expanding our benchmark.
+5.2.3. Alphabattle: Default
+In Table 2, we present the metrics for unimodal methods on the Alphabattle dataset.
+Interestingly, the supervised methods on both the MBD and Alphabattle datasets perform as well as, or better than, unsupervised methods.
+This may be attributed to the dataset size and the available labeled data.
+5.2.4. Multimodal matching task
+In this subsection, we present the results of our proposed multimodal matching benchmark, summarized in Table 4, which includes both MBD and Datafusion datasets.
 Detailed results for other modalities can be found in Appendix B, in Table 8.
-In this evaluation, we report recall@1, recall@50, and recall@100 for each modality, measured in both directions.
-For example, in the case of the transaction and geostream pair, we compute both Trx2Geo and Geo2Trx, allowing for an evaluation of alignment from both perspectives.
+We report Recall@1, Recall@50, and Recall@100 for each modality, measured in both directions.
+We also present the multimodal matching results for the DataFusion dataset.
+For example, in the case of the transaction and geostream pair for MBD, we compute both Trx2Geo and Geo2Trx, evaluating alignment from both perspectives.
 Our analysis reveals considerable variation in performance across different modality pairs.
-Specifically, dialogue data consistently exhibits weaker matching performance compared to other modalities, such as transactions and geostream, which demonstrate significantly stronger alignment.
+Specifically, dialogues consistently exhibit weaker matching performance than other modalities, such as transactions and geostream, demonstrating significantly stronger alignment.
 This disparity suggests potential limitations within the dialogue modality, indicating that it may offer less complementary or aligned information.
-Alternatively, the unique structure of dialogue data may necessitate more sophisticated or specialized techniques for effective integration with other modalities.
-These findings provide important insights for the advancement of multimodal matching methods, underscoring the need for further refinement.
-| Recall@1 | Recall@50 | Recall@100 | |
-| Trx2Geo | 0.006 0.0003 | 0.1967 0.002 | 0.303 0.004 |
-| Geo2Trx | 0.004 0.0003 | 0.1624 0.002 | 0.262 0.004 |
-5 Related Work
-5.1 Financial data
-The multitude of services and processes in banks generates a variety of data that can be considered as modalities.
-Early works such as [29, 26] use feature processing techniques that remove multimodal complexity and present data in tabular form.
-The Amex dataset[3] improves information content and complexity.
-Here, a wide range of different financial aggregates are presented as a sequence of historical slices.
-The development of deep learning methods has led to the ability to work with complexly structured data, such as sequences of events.
-Quite a few datasets [1, 5, 2], mostly monomodal, presented mainly at ML competitions.
-To work with such data, both supervised[7, 9, 36] and unsupervised methods[30, 8, 32] are used.
-Multimodal financial sequential data proposed in DataFusion2022 [4].
-There are two sequential modalities, transaction and web clickstream, and two downstream tasks: matching and education level prediction.
-However, this is an extremely small dataset of 17K clients, and no accurate baseline model is available.
-5.2 Other event sequence domains
-Temporal point process[28, 40] model streams of discrete events in continuous time by constructing a neurally multivariate point process.
-The authors use a large collection of datasets from different types of modalities: Media (Retweets[39], MemeTrack[24], Amazon[6], IPTV[25]), Medical (MIMIC-II[22]), Social(Stack Overflow[24], Linkedin[38]) and Financial (Transaction[15]) data.
-All datasets are independent, and each one is single-modal.
-EventStreamGPT [27] uses multimodal medical record datasets, MIMIC-IV[21].
-The authors propose a GPT-like approach for continuous-time event sequences.
-The structure of this dataset is close to the MDB.
-But it must be emphasized that financial data, unlike medical data, contains longer chains of events, more regular patterns, and individual transactions are less informative.
-These features may influence the quality of the methods used.
-5.3 Geostream and dialogues
-Geodata is used for various tasks.
-One of the uses of geo is a visualization of analytics on a map [17].
-However, geo is used here not as a separate modality but as additional tags to the mainstream of tweets.
-Mobile marketers[12] use geo-targeting for pricing and send personalized recommendations.
-In [35] geo hashes are used for user mobility detection and prediction.
-We encode our dialogue entries via a pretrained NLP Model.
-Previously, such models have already been used[18] for text anonymization tasks.
-Embeddings preserve the meaning of the text, which was shown in[34].
-Pre-trained text embeddings can capture text sentiment and improve text-to-speech models[19].
-6 Limitations
-The study used data from only one company, which may limit the generalizability of the results to other organizations.
-In addition, the data was subject to de-identification, which limits the possibility of using models trained on this dataset outside of it.
-Also, the data analysis results can not be generalized.
-In other words, based on this dataset, it is impossible to draw conclusions regarding specific regions and market characteristics or perform deep text analytics.
-However, within the dataset, the data is consistent, which allows us to draw correct conclusions about the performance of multimodal or unimodal methods for working with sequences.
-It is also worth noting that the study was conducted on a sample of clients of a certain segment who had the opportunity to purchase certain products and does not cover all possible groups of consumers.
-7 Ethics and Societal Implications
-Our dataset is designed to facilitate the development of algorithms for analyzing complex event data, with the goal of enhancing the accuracy and efficiency of predictive models.
-Our event data does not contain characteristics such as gender, age, or race, which can be used to discriminate against certain groups of people.
-This makes the models trained on our dataset more ethical.
-However, it is important to note that companies must comply with legal grounds and obtain customer consent before collecting and analyzing additional data.
-It is also worth mentioning the need to ensure the security of data storage and transmission to prevent leakage of confidential information.
-Failure to comply with these conditions may result in privacy violations.
-8 Conclusion and future work
-In this paper, we introduced the first large-scale publicly available multimodal banking dataset, MBD333https://huggingface.co/datasets/ai-lab/MBD, which contains anonymized sequential data (bank transactions, geo position, and technical support dialogues) for 1M bank clients.
-It can be used to create industrial multimodal models for event sequences to predict client needs.
-Using this dataset, we present the novel benchmark based on MBD with two practically important tasks: campaigning (purchase prediction for the next month) and matching different modalities of clients.
-Our experimental results demonstrate that even a simple fusion of multiple data sources improves the overall quality compared to single-modal baselines (Table 2).
-Also, we show that anonymization does not significantly impact the algorithms’ performance, so our dataset is appropriate for choosing the best model that can be implemented in production scenarios.
-As a result, our work opens new horizons for developing practically relevant large-scale multimodal algorithms for event sequences.
-In the future, expanding the range of tasks with our dataset is possible.
-For example, it is possible to estimate the multimodal techniques of the next-event prediction.
-Moreover, or baseline methods are based on late fusion, so it is very important to study more sophisticated models that consider possible mutual influence between modalities.
-Appendix A Dataset Statistics
+Alternatively, the unique structure of dialogue data may necessitate more sophisticated techniques for effective integration with other modalities.
+| Dataset | Modalities | Recall@1 | Recall@50 | Recall@100 |
+| MBD | Trx2Geo | 0.006 0.0003 | 0.196 0.002 | 0.303 0.004 |
+| Geo2Trx | 0.004 0.0003 | 0.162 0.002 | 0.262 0.004 | |
+| Datafusion | Trx2Click | 0.002 0.0010 | 0.063 0.004 | 0.120 0.009 |
+| Click2Trx | 0.001 0.0007 | 0.070 0.005 | 0.115 0.008 |
+| Dataset | Modalities | Blending | Late Fusion | Early Fusion | |
+| TabGPT | TabGPT | Supervised RNN | CrossTransformer | ||
+| MBD | Trx + Geo | 0.804 0.001 | 0.800 0.001 | 0.819 0.001 | 0.815 0.001 |
+| Trx + Dialog | 0.742 0.001 | 0.810 0.002 | 0.821 0.0006 | 0.821 0.002 | |
+| Datafusion | Trx + Click | 0.756 0.013 | 0.766 0.011 | 0.703 0.008 | 0.735 0.010 |
+5.3. Comparison of Multimodal Fusion Methods
+In this subsection, we evaluate the performance of multimodal fusion techniques across multiple datasets.
+For this analysis, we select the best-performing models for Blending and Late Fusion to compare against Early Fusion.
+As shown in Table 5, Late Fusion with a Supervised RNN consistently delivers superior results on the MBD dataset, demonstrating its robustness in effectively integrating multimodal data.
+Early Fusion, implemented via the cross-attention from CrossTransformer (Zhang and Yan, 2023), achieves competitive performance but slightly lags behind Late Fusion, while Blending exhibits considerably weaker results.
+On the Datafusion dataset, Late Fusion with TabGPT is the most effective method, with Early Fusion also performing well and outperforming Late Fusion with a Supervised RNN.
+Blending, in contrast, consistently yields the lowest performance across all modalities.
+These findings highlight the effectiveness and reliability of Late Fusion for multimodal data integration while identifying Early Fusion as a promising direction for further research and optimization.
+5.4. LLMs for multimodal banking data
+Large Language Models, or LLMs, are now strong tools for handling various types of tabular data (Ruan et al., 2024).
+This includes sequential tabular data (Zhang et al., 2023).
+We used popular models LLAMA3.2 (Meta, 2024) and QWEN2.5 (Yang et al., 2024).
+These two are publicly available.
+We also used two in-house LLMs, Gigar4 and Gigar4 SFT (Supervised Fine-Tuning).
+The last two models are Russian language-specific language models.
+Gigar4 SFT is a version that has been fine-tuned on several math datasets.
+The fine-tuning of math datasets enhances their ability to process and predict numerical data accurately.
+We took 3B parameter versions of the models to keep inference time low while being able to process context up to 8192 tokens on a single GPU.
+Experiments with LLMs are computationally expensive, so we used the MBD-small subset described in Subsection 3.1.
+Inference on this dataset using 3B parameters LLMs takes 12 hours on a single NVIDIA A100 GPU.
+Our results (Table 6) show that in transactions, relatively small language models can match the performance of specialized event-sequence models, even without extra tuning.
+At the same time, there is room for improvement to better account for data-specific characteristics.
+Geo hashes exhibit extreme diversity in values (44K unique anonymized values even for level-4 hashes), which makes it challenging for models to generalize without fine-tuning.
+Additionally, adding modalities increases the required context length, which is already a challenge for LLMs, given the long event sequences and the way they are encoded into text.
+| Model | Trx | Trx+Geo | Geo | |
+| Gigar4 | 0.7730.007 | 0.7370.006 | 0.5480.004 | |
+| LLM | Gigar4 SFT | 0.7780.005 | 0.7350.014 | 0.5490.010 |
+| LLAMA3.2 | 0.7730.006 | 0.7350.011 | 0.5340.016 | |
+| QWEN2.5 | 0.7570.005 | 0.7270.007 | 0.5390.008 | |
+| CoLES | 0.7850.006 | 0.7820.005 | 0.5680.007 | |
+| TabGPT | 0.7790.007 | 0.7820.002 | 0.5940.016 | |
+| Baselines | TabBERT | 0.7810.003 | 0.7890.002 | 0.6010.013 |
+| Aggregation | 0.7810.009 | 0.7830.008 | 0.5300.006 | |
+| Supervised RNN | 0.7450.004 | 0.7420.006 | 0.5320.013 |
+6. Conclusion and Future Work
+In this paper, we introduce MBD, the first large-scale, publicly available multimodal banking dataset (Table 1), comprising sequential data from over 2 million clients, including transactions, geo-locations, and support dialogues.
+MBD ensures compliance with data protection regulations, enabling research on financial data while maintaining privacy.
+Our findings show that anonymization has minimal impact on algorithm performance (Table 3, Fig.
+2), making the dataset suitable for real-world deployment.
+Additionally, excluding sensitive attributes (e.g., gender, age, race) mitigates bias, fostering ethical AI development.
+MBD, along with Datafusion and Alphabattle, establishes a benchmark for key downstream tasks, facilitating the development of scalable multimodal and unimodal algorithms for event sequence prediction.
+We implemented the state-of-the-art models in banking domains, demonstrating that even basic multimodal fusion outperforms single-modal baselines (Table 3).
+Further, Late and Early Fusion experiments (Table 5) highlight the potential for improving model performance by effectively capturing cross-modal interactions.
+We also assess LLMs in unimodal and multimodal settings, comparing them with domain-specific approaches (Table 6).
+Given MBD’s scale and practical relevance, even modest performance gains can yield significant financial benefits.
+Future iterations will incorporate additional data sources (e.g., clickstreams, enriched dialogue features) and expand downstream tasks to financial product recommendations, customer churn prediction, fraud detection, and credit risk assessment.
+Extending the dataset’s time range will further enable longitudinal studies and trend analysis
+Appendix A Dataset Statistics and data samples
+In this section, we present data samples from various modalities and a histogram of sequence lengths.
+Table 7 provides a list of attributes for the transactional modalities, while Figure 3 illustrates the distribution of sequence lengths for these modalities.
+Additionally, Figure 6 shows a sample of geographical data, and Figure 7 displays the distribution of sequence lengths for geostream modalities.
+For dialogue data, a sample is presented in Figure 4, with the distribution of sequence lengths for dialogue events shown in Figure 5.
+Figure 8 demonstrates a sample of data for product purchases.
+Here, we observe that geographical and transactional modalities have long tails in the histogram of distributions.
 | Features | Number of categories | Description |
@@
 | src type32 | 88 | sender field subtype 3 |
+The histogram of the number of clients with a certain length of transaction history.
+Sample data of dialogues modality.
+The histogram of the number of clients with a certain number of dialogues.
+Sample data of geostream modality.
+The histogram of the number of clients with a certain length of geostream.
+Example data of client’s purchases.
+Median amount per client for private transaction data.
+Mean sequence length per client for private transaction data.
 Appendix B Detailed experimental results
+B.1. Multimodal matching
+For the multimodal matching task Table 8, it is observed that the dialogue modality exhibits poor compatibility when combined with other modalities.
+| Recall@1 | Recall@50 | Recall@100 | |
+| Trx2Geo | 0.006 0.0003 | 0.1967 0.002 | 0.303 0.004 |
+| Geo2Trx | 0.004 0.0003 | 0.1624 0.002 | 0.262 0.004 |
+| Trx2Dial | 0.00004 0.00001 | 0.0001 0.00005 | 0.003 0.0001 |
+| Dial2Trx | 0.00003 0.00001 | 0.001 0.00005 | 0.003 0.0001 |
+| Dial2Geo | 0.00002 0.000001 | 0.0006 0.00001 | 0.001 0.0001 |
+| Geo2Dial | 0.00002 0.000001 | 0.0005 0.00001 | 0.001 0.0001 |
+B.2. Handling class imbalance
+We conducted experiments on the MBD dataset, focusing on transaction modalities, and explored various techniques, including random undersampling, oversampling, and class balancing within gradient boosting for CoLES embeddings.
+Additionally, stratified batching was applied to better align the target distribution in Supervised RNN.
+The results of these experiments are summarized in Table 9.
+| Method | Description | ROC-AUC |
+| Supervised RNN | Baseline | 0.819 |
+| Supervised RNN (stratified batching) | Balanced target distribution | 0.819 |
+| CoLES | Baseline embeddings | 0.773 |
+| CoLES (undersampling) | Random undersampling applied | 0.772 |
+| CoLES (oversampling) | Random oversampling applied | 0.774 |
+| CoLES (class balanced) | Class balancing in gradient boosting | 0.774 |
+Here, stratified batching consistently maintains a high ROC-AUC of 0.819 for Supervised RNN, demonstrating its robustness to label imbalance.
+For CoLES embeddings, balancing techniques result in minimal variations in performance, with ROC-AUC values ranging from 0.772 to 0.774.
+This indicates that CoLES embeddings exhibit limited sensitivity to label imbalance, highlighting the potential need for more advanced balancing strategies to achieve further improvements.
+B.3. Campaining results on the private dataset and MBD
+The experimental results presented in Tables 12 and 13 underscore the pivotal importance of modality fusion in enhancing model performance.
+Here, integrating dialogue data with transaction data leads to a significant performance boost.
+For instance, augmenting TrxTabGPT with DialogLast results in a 1.8% improvement in the mean metric for blending and a 3.4% enhancement in fusion.
+The most substantial performance gains are observed when all modalities are combined.
+Specifically, the integration of TrxTabGPT, GeoTabGPT, and DialogLast yields a 2% improvement over unimodal transaction models and a 3.2% improvement in late fusion, highlighting the synergistic benefits of incorporating dialogue and geographical data into transaction-based models.
+These findings provide robust evidence supporting the effectiveness of multimodal integration.
+The inclusion of dialogue and geographical data significantly boosts the performance of models centered on transaction data.
+Moreover, this trend observed in public datasets is consistently replicated in proprietary data, as shown in Tables 10 and 11.
+These results also prove the low impact of our anonymization procedure.
+Indeed, the ranking of methods remains consistent despite differences in absolute metric values.
+As shown in Table 11, TrxTabGPT (ROC-AUC 0.796) and TrxAggregation (ROC-AUC 0.780) achieve the best performance on the private dataset.
+Similarly, TrxTabGPT leads on the public dataset with an ROC-AUC of 0.802 (Table 13).
+Incorporating geolocation and dialogue modalities further improves results, with DialogLast+TrxTabGPT+GeoTabGPT attaining the highest ROC-AUC on both datasets: 0.802 on the private dataset (Table 11) and 0.808 on the public dataset (Table 13).
+Overall, multimodal approaches utilizing TabGPT or Aggregation demonstrate superior performance, with Late Fusion consistently outperforming Blending across private and public datasets (compare results in Table 10 and Table 12 to those in Table 11 and Table 13).
 | Methods | mean | target_1 | target_2 | target_3 | target_4 |
@@
 | TrxTabBERT+GeoTabBERT | 0.764 0.004 | 0.716 0.006 | 0.823 0.004 | 0.737 0.006 | 0.780 0.005 |
-| Recall@1 | Recall@50 | Recall@100 | |
-| Trx2Geo | 0.006 0.0003 | 0.1967 0.002 | 0.303 0.004 |
-| Geo2Trx | 0.004 0.0003 | 0.1624 0.002 | 0.262 0.004 |
-| Trx2Dial | 0.00004 0.00001 | 0.0001 0.00005 | 0.003 0.0001 |
-| Dial2Trx | 0.00003 0.00001 | 0.001 0.00005 | 0.003 0.0001 |
-| Dial2Geo | 0.00002 0.000001 | 0.0006 0.00001 | 0.001 0.0001 |
-| Geo2Dial | 0.00002 0.000001 | 0.0005 0.00001 | 0.001 0.0001 |